### v3

# HBS - Hardware Build System: A Tcl-based, minimal common abstraction approach for build system for hardware designs

Abstract.
Build systems become an indispensable part of the software implementation and deployment process. New programming languages are released with the build system integrated into the language tools, for example, Go, Rust, or Zig. However, in the hardware description domain, no official build systems have been released with the predominant Hardware Description Languages (HDL) such as VHDL or SystemVerilog. Moreover, hardware design projects are often multilanguage.
The paper proposes a new build system for the hardware description domain. The system is called the Hardware Build System (HBS). The main goals of the system include simplicity, readability, a minimal number of dependencies, and ease of integration with the existing Electronic Design Automation (EDA) tools. The system proposes a novel, minimal common abstraction approach, whose particular implications are described in the article. All the core functionalities are implemented in Tcl. Only the EDA tool’s independent features, such as dependency graph generation, are implemented in a Python wrapper.
Automation, Build System, Hardware Synthesis, Project Maintenance, Testbench Runner
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Hardware Software tools for EDA
1. Introduction
Implementation of any complex software or hardware description design involves not only code writing. Equally important is defining metadata used for configuration, dependency management, code generation, testing, and deployment purposes. The metadata is read by software tools, which, in turn, are responsible for generating final binary or bitstream files, or for carrying out extra tasks such as unit tests. Such programs are often referred to as build systems. Some build systems are split into multiple smaller programs, and some have a monolithic structure. An introduction to the build system concept and principles can be found in (Smith, 2011; MOKHOV et al., 2020).
In the software domain, there are multiple well-established build systems, such as GNU Make (Free Software Foundation, 2025), CMake (CMake, 2025), Ninja (Martin, 2025), or Bazel (Google, 2025). Some of them were implemented in the early days of programming. For example, GNU Make was first released in 1988 as part of the GNU Project. It was implemented to provide a free software alternative to Unix Make, which was first implemented in 1976 by Stuart Feldman at Bell Labs. Modern programming languages, such as Go, Rust, or Zig, come with build systems included in the language toolchains. The Zig build system can even be used as a build system for C/C++ projects.
In the hardware domain, the situation looks different. The earliest Electronic Design Automation (EDA) tools were implemented as proprietary programs. EDA vendors did not (and still do not) cooperate to create a standard (or at least similar) user interface for tool interaction. The main common point between various EDA tools is using the Tcl language as the interpreter for interaction with the tools. However, the tools provide custom commands and parameters even for primary, common tasks, such as adding hardware description files to a project design. For example, to add a VHDL file in AMD Vivado, one can use specific read_vhdl <path_to_vhdl_file> or generic add_files <path_to_vhdl_file> commands. To add a VHDL file in Intel Quartus, one has to use set_global_assignment -name VHDL_FILE <path_to_vhdl_file> command. EDA tools have abundant differences, making implementing a build system for hardware projects harder than for software. Despite this, there were (and still are) attempts to implement build systems that are more suitable for hardware projects.
The paper proposes a new build system for a hardware description domain. The system is Tcl-based, which makes it easy to integrate with all existing EDA tools already utilizing Tcl language as a user-tool interface. A novel, minimal common abstraction approach differentiates the implemented system from the existing build systems targeting the hardware domain.
The structure of the article is as follows. Section 2 describes related work. Section 3 discusses the requirements and their reasons imposed on the implemented build system. Section 4 introduces the internal architecture of the implemented build system. Section 5 describes commands provided by the implemented build system. Section 6 evaluates the proposed solution in terms of other build systems. Section 7 summarizes the work and highlights novel aspects of the implemented build system.
2. Related work
There already exist at least more than ten build systems targeting the hardware domain. Some of them are vendor tool specific, and some try to be generic, entirely abstracting away the tool used for synthesis or implementation. By looking at the architecture (analyzing how a tool internally works), existing hardware build systems can be classified into two approaches:
  1. (1)
    direct Tcl approach,
  2. (2)
    indirect abstract approach.
2.1. Direct Tcl approach
In the Tcl approach, the user directly utilizes the Tcl interface. The idea seems obvious. Since EDA tools already utilize Tcl as a user interface, a build system should also provide users with a Tcl interface. Usually, the user does not directly call EDA tool specific commads. Tcl approch build systems implement some wrapper commands that the user should use. Representatives of the direct Tcl approach are, for example, fpga-vbs (Braun, 2025), vextproj (Zabołotny, 2025, 2016), and OSVVM-Scripts (Lewis, 2025).
Being able to easily execute arbitrary, custom EDA tool commands is the most significant advantage of the direct Tcl approach. As a Tcl interpreter embedded into an EDA tool executes the build system code, a user can call custom commands in almost any place.
The direct Tcl approach is not free of drawbacks. Some of them are inherent, and some result from the fact that direct Tcl build systems development was stopped at some stage by their maintainers.
The first inherent drawback is that Tcl is considered an unfriendly language compared to, for example, Python. The argument is often raised by junior hardware engineers and has a very subjective nature. However, Tcl was chosen as the EDA scripting language because it was very popular when EDA tools were developed (Chen et al., 2001). The second inherent drawback is that Tcl is not a popular language nowadays. Although there are no physical obstacles to change this trend, it is doubtful that suddenly more developers start using Tcl instead, for example, Python as an interpreted language. This is why this drawback is also treated as inherent.
The second group of direct Tcl build systems drawbacks results from stopping the development at a certain stage. There is no direct Tcl build system that can be used to examine the limits and bottlenecks of this idea. For example, most of the direct Tcl build systems support only one EDA tool. OSVVM-Scripts is an exception. However, it can be used only for simulation. It does not support any synthesis or implementation tools. More example drawbacks common to the direct Tcl build systems include:
  1. (1)
    No dependency graph generation capability.
  2. (2)
    Lack or limited capability of automatic testbench targets detection.
  3. (3)
    Lack of automatic parallel testbench running.
2.2. Indirect abstract approach
The indirect abstract approach is much more popular than the direct Tcl approach. In this approach, the user declares various properties and configurations instead of directly writing Tcl code or calling EDA tools Tcl commands.
The most popular hardware build system utlizing the indirect abstract approach is probably FuseSoc (Kindgren, 2019, 2025b). Other popular build systems include SiliconCompiler (Olofsson et al., 2022; sil, 2025), HAMMER (Liew et al., 2022; ham, 2025), Hog (Biesuz et al., 2021; Gonnella et al., 2025), hdlmake (hdl, 2025), orbit (Ruskin, 2025), bender (ben, 2025), bazel_rules_hdl (baz, 2025), flgen (Ishitani, 2025), and Blockwork (blo, 2025).
The language and syntax used for declarations vary between different build systems. Some use simple data serialization formats such as YAML (for example, FuseSoc or bender) or TOML (for example, orbit). Some come with a domain-specific language (for example, flgen), or use Python (for example, HAMMER) and Makefiles (for example, Hdlmake). Finally, some use a mix of more than one (for example, Hog). Especially using Python as a declarative language is powerful, as users can use all functionalities that come with the language.
The most significant advantage of the indirect abstract approach is the low entry threshold. The syntax of languages such as YAML, TOML, or Python is simple, which leads to high legibility. Even junior engineers without extra training can understand it. Listing 1 presents a very basic core definition in the FuseSoc build system. The first line specifies the version of the description format. The “name” key specifies the vendor, library, name, and version (VLNV) identifier specified in the IP-XACT User Guide (Accellera IP-XACT Working Group, 2018). The remaining part is relatively self-explanatory.
Listing 1: Example edge detector core definition in FuseSoc build system.
CAPI=2:
name: vendor:library:edge_detector:1.0
filesets:
src:
files:
- src/edge_detector.vhd
file_type: vhdlSource-2008
logical_name: hdl_library_name
targets:
default:
filesets:
- src
The indirect abstract approach is based on the concept that the interface to EDA tools is hidden under some common abstraction layer. Whether this is an advantage or a disadvantage is debatable. Providing a common abstract layer means that users do not have to learn the interface of multiple EDA tools. Moreover, in theory, changing the EDA tool for compilation process should be limited to changing solely the EDA tool declaration. However, this is somewhat a utopian view. As practice shows the common abstraction layer leads to the following common difficulties:
  1. (1)
    Missing feature implementation for some EDA tools. Let us suppose the abstraction layer declares feature F, and the build system claims support for tools T1 and T2. Although T1 might provide the implementation for F, T2 might be missing implementation for this particular feature. In such a case, the user must add support for this feature to the build system. This task might not be trivial. The user must get to know the internals of the build system, prepare a pull request, wait for a review, and potentially reiterate the implementation. Yet another approach is to modify automatically generated build scripts before a compilation. However, this is against the whole idea of build systems. A direct Tcl approach is free of these difficulties because the user can directly insert custom Tcl commands. There are no intermediate abstraction layers.
  2. (2)
    Simple tasks, such as scoping constraints to a particular module, sometimes require verbose solutions. The solution is not only more verbose but also more complex to understand. What can be solved in one or two lines of Tcl code requires more than one file. An example is discussed in subsection 6.3.
All indirect abstract hardware build systems allow “injecting” arbitrary custom Tcl commands into the automatically generated scripts. However, these commands can not be inserted into arbitrary places. Moreover, inserting a single-line Tcl command usually requires more than one line of code or even more than one file. This can be treated as a proof that the indirect abstract approach is not a golden bullet. The idea of a common abstraction layer has significant drawbacks that no one yet discovered how to fix.
Making a truly abstract hardware build system is highlighy unliekly in the nearest feature. The main problem with the common abstraction layer is the fact that EDA tools were implemented before such a layer was proposed. Declaring a common interface for already existing tools is simply much harder task than implementing tools conforming to some formal interface.
The HBS described in the paper proposes a novel approach. It tries to incorporate the best of the direct Tcl approach and indirect abstract approach. It implements a common abstraction layer. However, the layer is:
  1. (1)
    minimal,
  2. (2)
    limited to the primary common features,
  3. (3)
    implemented in Tcl.
3. Hardware Build System requirements
3.1. Simplicity and ease of use
The overhead of build system implementation and maintenance is often underrated (Nejati et al., 2024; Adams, 2009). McIntosh in (McIntosh, 2011) reports that the build system can account for up to 31% of the code files in a project. Moreover, up to 27% of development tasks that change the source code also require build system maintenance. The study was carried out on software projects, not on hardware designs. However, due to the complexity of hardware design, it seems reasonable to assume comparable shares.
Nejati et al. in (Nejati et al., 2023) report that the complexity of build system specifications is one of the top three factors complicating the review of build specifications.
The primary goal of the hardware build system proposed in the paper is simplicity. Contemporary hardware build systems utilizing the indirect abstract approach offer simplicity only at the initial stage. However, as the project complexity grows, so does the complexity of the build system utilization.
3.2. Minimal user interface
Xia et al. in (Xia et al., 2013) report that even up to 21.35% of the bugs in software build systems belong to the external interface category. This means that one-fifth of build system bugs results from incorrect usage of the user interface or different behavior of the build system with different tools or platforms.
Fixing build system external interface bugs requires applying patches in the build system itself. Such a task is relatively time-consuming and might not be simple to implement. Build system user must report the problem to the build system developers or fix the bug himself. However, this implies that the build system user is distracted from his own work. Moreover, the fix is not applied immediately. This means that the user must also find a temporary workaround. This, in turn, leads to a technical debt (Besker et al., 2019, 2018).
A minimal user interface has one more advantage. It allows developers to customize build systems to the project needs. Each project is in some way unique and might require particular approach for build problems solving. Moreover, projects usually evolve, for example, due to the customer demand. Dynamic tailoring to the project requirements is a desired feature of modern build systems (Maudoux and Mens, 2018). Similar observations were made by Neitsh et al. in (Neitsch et al., 2012). Authors report that concerns from the application and implementation domains may “leak” into the build model. The build system should be adaptable enough to provide ways to respond to these concerns. Too rigid user interfaces limit such capabilities.
3.3. Low number of dependencies
Xiao et al. in (Xiao et al., 2022) report that dependency management is one of the top three factors of technical debt introduction in build systems. Such a finding is not surprising. Dependencies often break in unexpected moments. In such a case, long-term quality is sacrificed for the short-term goals. Getting a project to compile successfully is more important than fixing broken build system dependencies.
One of the goals of the HBS proposed in the paper is to minimize the probability of user distraction caused by broken dependencies. This implies that the number of dependencies should be kept as low as possible.
Moreover, a low number of dependencies potentially makes the build system more appealing for use in military designs. This is because any external dependency must be verified against potential vulnerabilities.
3.4. Ease of calling arbitrary programs in arbitrary places
The ease of calling arbitrary programs in arbitrary places results from two use cases.
The first use case originates from the complexity of designing hardware systems. To reduce manual work, engineers often automatically generate parts of the design. For example, register files are automatically generated from SystemRDL (Accellera SYSTEMS INITIATIVE, 2025). In such a case, there must be a way to call an external program generating hardware description of the register components. Existing indirect abstract hardware build systems allow calling external programs. However, they often require users to define a transitional so-called generator, which increases complexity and reduces legibility. Moreover, they usually do not allow to call external programs in arbitrary places.
The second use case includes support for EDA tools without the Tcl interface. Such tools are often open source simulators, for example, nvc (N., 2025), or verilator (W., 2025). These tools have interfaces similar to the standard software compiler. The user simply calls programs with proper arguments and proper order in a shell. Being able to call arbitrary programs in arbitrary places is required to support this kind of tools without an extra script or Makefile generation.
3.5. Independce from version control system
Some hardware build systems are tightly coupled with a particular version control system. For example, Hog is intended to be used only with git. However, build systems and version control systems are conceptually orthogonal entities. Imposing a particular version control system limits user choice. Moreover, it potentially makes incorporating a new build system for some existing projects impossible, as they might, for example, utilize subversion instead of git.
4. Hardware Build System architecture
This section describes the architecture of the implemented HBS. The system is available on GitHub (M., 2025) with the BSD 3-Clause License.
It was decided that HBS will be implemented using Tcl language because of the following reasons:
  1. (1)
    Implementing a hardware build system in Tcl allows the execution of build system code during the EDA tool flow. This, in turn, gives direct access to all EDA tool custom commands. Moreover, these custom commands can be evaluated in arbitrary places. This aspect satisfies requirements 3.1 and 3.4. The complexity of existing indirect abstract build systems for less common tasks results from executing build system code before EDA tool flow. The build system only prepares scripts that are later run by EDA tools.
  2. (2)
    If the EDA tool provides the Tcl interface, then the Tcl shell is provided by the EDA tool vendor. The shell is installed during the installation of vendor tools. This implies that, in some cases, the build system user does not even have to install additional programs. This aspect satisfies requirement 3.3.
  3. (3)
    Executing arbitrary programs in arbitrary places in Tcl is very simple. There is a dedicated exec command for invoking subprocesses. If executing a subprocess requires prior dynamic arguments evaluation, the exec command call must be prepended with the eval command. Even in Python, invoking a subprocess is not so straighforward. This Tcl feature satisfies requirement 3.4.
One of the most important things while designing the HBS was the separation of common build process operations that would constitute a common abstraction layer over EDA tools. At the end, it was decided that the following actions should constitute the common abstraction layer:
  1. (1)
    Target device setting - some EDA tools use the term “part” instead of “device,” for example, Vivado. Simulation EDA tools do not require a device setting. However, all synthesis EDA tools require information on the target device. This is why setting the device became part of the common abstraction layer.
  2. (2)
    File addition - this includes support for adding files of all formats supported by a given EDA tool.
  3. (3)
    Library setting - setting HDL file library.
  4. (4)
    HDL standard setting - setting HDL file standard revision. The build system has to manage this because some tools can not handle analyzing different design units using different standard revisions, for example, nvc. In such a case, the build system must decide what the common standard revision shall be used for analyzing all HDL files.
  5. (5)
    Dependency specification - this is the core feature of any build system.
  6. (6)
    Generics/parameters setting - configuring parametric designs must be an inherent feature of any hardware build system.
  7. (7)
    Design top module setting - all EDA tools carrying out simulation or synthesis requires information on the top module.
  8. (8)
    Code generation - a hardware build system must provide a simple way for automatic arbitrary code generation. This is further explained in subsection 4.12.
4.1. General structure
The HBS is implemented in two files hbs.tcl and hbs. The first is implemented in Tcl, and the second in Python. The hbs.tcl file implements all the core features related directly to the interaction with the EDA tools. The hbs.tcl provides the following functions:
  1. (1)
    Dumping information about detected cores in JSON format.
  2. (2)
    Listing cores found in .hbs files.
  3. (3)
    Listing targets for a given core.
  4. (4)
    Running target provided as a command line argument.
The hbs is a wrapper for the hbs.tcl and serves the following additional functions:
  1. (1)
    Showing documentation for hbs Tcl symbols.
  2. (2)
    Generating dependency graph.
  3. (3)
    Listing testbench targets.
  4. (4)
    Running testbench targets.
By default, the user calls the hbs program. However, if none of the additional functions are required, the user can call the hbs.tcl directly. In such a case, the whole build system is limited to a single file. Check Table 1 to see how it compares to the size of existing hardware build systems.
4.2. Cores and cores detection
When the user executes hbs (or hbs.tcl), all directories, starting from the working directory, are recursively scanned to discover all files with the .hbs extension (symbolic links are also scanned). Files with the .hbs extension are regular Tcl files that are sourced by the hbs.tcl script. However, before sourcing .hbs files, the file list is sorted so that scripts with shorter path depth are sourced as the first ones. For example, let us assume the following three .hbs files were found:
  • a/b/c/foo.hbs,
  • d/bar.hbs,
  • e/f/zaz.hbs.
Then, they would be sourced in the following order:
  1. (1)
    d/bar.hbs,
  2. (2)
    e/f/zaz.hbs,
  3. (3)
    a/b/c/foo.hbs.
Such an approach allows controlling when custom symbols (Tcl variables and procedures) are ready to use. For example, if the user has a custom procedure used in multiple .hbs files, then the user can create separate utils.hbs file contaning utility procedures, and place it in the the project root directory. Within .hbs files, the user usually defines cores and targets, although the user is free to have any valid Tcl code in .hbs files.
Listing 2 presents a very basic flip-flop core definition. The flip-flop core has a single target named src. The core consists of a single VHDL file.
Listing 2: Example flip-flop core definition in HBS.
namespace eval flip-flop {
proc src {} {
hbs::AddFile flip-flop.vhd
hbs::Register
To register a core, the user must explicitly call hbs::Register procedure at the end of the core namespace. Such a mechanism helps to distinguish regular Tcl namespaces from Tcl namespaces representing core definitions. If the user forgets to register a core, the build system gives a potential hint. An example is presented in Listing 3.
Listing 3: An HBS error message with a hint on core registration.
[user@host tmp]\$ hbs run lib::core::tb
core lib::core not found, maybe the core is not registered (hsb::Register)
Each core is identified by its unique path. The core path is equivalent to the namespace path in which hbs::Register is called. Using the namespace path as the core path gives the following possibilities:
  1. (1)
    The user can easily stick to the VLNV identifiers if required. This is presented in Listing 4. In this case, the flip-flop core path is vendor::library::flip-flop::1.0.
  2. (2)
    The user can define arbitrary deep core paths (limited by the Tcl shell). This is presented in Listing 5. In this case, the core path consists of 7 parts.
  3. (3)
    The user can nest namespaces to imitate the structure of libraries and packages. This is presented in Listing 6. Three flip-flop cores are defined in the snippet. Listing 7 presents output for listing flip-flop cores.
Listing 4: Example flip-flop core definition with VLNV core path.
namespace eval vendor::library::flip-flop::1.0 {
proc src {} {
hbs::AddFile flip-flop.vhd
hbs::Register
Listing 5: Example flip-flop core definition with 7 core path parts.
namespace eval a::b::c::d::e::f::flip-flop {
proc src {} {
hbs::AddFile flip-flop.vhd
hbs::Register
Listing 6: An example .hbs file with three flip-flops core definitions
namespace eval lib {
namespace eval pkg1 {
namespace eval d-flip-flop {
proc src {} {
hbs::AddFile d-flip-flop.vhd
hbs::Register
namespace eval t-flip-flop {
proc src {} {
hbs::AddFile t-flip-flop.vhd
hbs::Register
namespace eval pkg2 {
namespace eval jk-flip-flop {
proc src {} {
hbs::AddFile jk-flip-flop.vhd
hbs::Register
Listing 7: Listing flip-flop cores output.
[user@host tmp]\$ hbs list-cores
lib::pkg1::d-flip-flop
lib::pkg1::t-flip-flop
lib::pkg2::jk-flip-flop
4.3. Targets and targets detection
HBS automatically detects targets. Targets are all Tcl procedures defined in the scope of core namespaces (namespaces with a call to hbs::Register). However, to allow users to define custom utility procedures within cores, procedures with names starting with the floor character (_) are not treated as core targets. Listing 8 presents an example edge detector core definition. The core path is vhdl::simple::edge-detector. The core has three targets: src, tb-sync, tb-comb, and one utility procedure _tb. The _tb procedure was defined to share calls common for testbench targets tb-sync and tb-comb. Moreover, all target procedures are also regular Tcl procedures. Such an approach allows for calling them in arbitrary places. The _tb procedure calls the src procedure because the edge_detector.vhd file is required for running testbench targets.
All targets are represented by a unique target path. Target path consists of the core path and target name. For example, the src target of the edge detector has the following path vhdl::simple::edge-detector::src.
Listing 8: Example edge detector core definition.
namespace eval vhdl::simple::edge-detector {
proc src {} {
hbs::SetLib "simple"
hbs::AddFile src/edge_detector.vhd
proc _tb {top} {
hbs::SetTool "ghdl"
hbs::SetTop \$top
src
hbs::SetLib ""
proc tb-sync {} {
_tb "tb_edge_detector_sync"
hbs::AddFile tb/tb_sync.vhd
hbs::Run
proc tb-comb {} {
_tb "tb_edge_detector_comb"
hbs::AddFile tb/tb_comb.vhd
hbs::Run
hbs::Register
4.4. Testbench targets
HBS is capable of automatically detecting testbench targets. Testbench targets are targets which names:
  1. (1)
    start with tb- or tb_ prefix,
  2. (2)
    end with -tb or _tb suffix,
  3. (3)
    equal tb.
For example, for the edge detector from Listing 8, the hbs program detects testbench targets presented in Listing 9.
Listing 9: Testbench targets listed for example edge detector core.
[user@host tmp]\$ hbs list-tb
vhdl::simple::edge-detector::tb-comb
vhdl::simple::edge-detector::tb-sync
4.5. Running targets
HBS allows running any target of registered cores. Even if the target itself has nothing to do with the hardware design. For example, running target print from Listing 10 results with the output shown in Listing 11.
Listing 10: Example print target not related to hardware build.
namespace eval core {
proc print {} {
puts "Hello!"
hbs::Register
Listing 11: Running print target result.
[user@host tmp]\$ hbs run core::print
Hello!
However, in most cases, the user wants to run a target related to the flow of the set EDA tool. In such a case, instead of manually calling all of the required tool commands, the user can call the hbs::Run procedure in the core target procedure. The hbs::Run procedure has an optional argument accepting the stage after which the tool flow should stop. This is further described in subsection 4.9. After hbs::Run returns, the user can continue processing. For example, the user can run scripts analyzing code coverage or preparing additional reports.
4.6. Target parameters
As core targets are just Tcl procedures, they can have parameters. Moreover, parameters can have optional default values. Additionally, HBS allows to provide command line arguments to the run target. This is a very convenient feature in build systems. Listing 12 presents a very simplified example.
Listing 12: Example target with an optional parameter.
namespace eval core {
proc target {{stage "bitstream"}} {
puts "Running until \$stage"
# hbs::Run commented out because this is just an example.
#hbs::Run \$stage
hbs::Register
The core does not build any hardware design. However, the example shows how the build stage can be passed from the command line to an EDA tool. Listing 13 presents output from running the target with different stage parameter values.
Listing 13: Running print target result.
[user@host tmp]\$ hbs run core::target
Running until bitstream
[user@host tmp]\$ hbs run core::target synthesis
Running until synthesis
Another practical example of target parameters usage is setting the simulator for testbench target from the command line or changing the top-level module. What target parameters are used for is limited only by the user’s imagination, and Tcl semantics.
4.7. Target context
An engineer implementing a given core controls the dependencies of the core. However, the core author does not control who will use the core and how. As targets are regular Tcl procedures, there is a need for a mechanism allowing the core author to evaluate the target procedure in the invariant environment. Such a mechanism in HBS is called the target context.
The target context assures that the following variables are not affected by dependee or dependency target execution:
  1. (1)
    HDL library,
  2. (2)
    HDL standard,
  3. (3)
    top module name,
  4. (4)
    arguments prefix,
  5. (5)
    arguments suffix,
  6. (6)
    the core path,
  7. (7)
    the target name,
  8. (8)
    the target path.
Listing 14 presents an example of the target context mechanism.
Listing 14: Target context example.
namespace eval pkg {
namespace eval foo {
proc src-foo {} {
hbs::SetLib "lib-foo"
hbs::AddDep pkg::bar::src-bar
puts "foo lib: \$hbs::Lib"
puts "foo core: \$hbs::ThisCorePath"
puts "foo target: \$hbs::ThisTarget"
hbs::Register
namespace eval bar {
proc src-bar {} {
hbs::SetLib "lib-bar"
puts "bar lib: \$hbs::Lib"
puts "bar core: \$hbs::ThisCorePath"
puts "bar target: \$hbs::ThisTarget"
hbs::Register
Listing 15 presents the output from running the pkg::foo:src-foo target. As can be seen, setting library in a target of one core, does not affect library in the target of another core.
Listing 15: Output from target context example.
[user@host tmp] hbs run pkg::foo::src-foo
bar lib: lib-bar
bar core: pkg::bar
bar target: src-bar
foo lib: lib-foo
foo core: pkg::foo
foo target: src-foo
An author of a core can not easily by mistake break execution of another core target procedure. The target context provides the isolation of target procedure evaluation environments.
4.8. Target dependencies
In HBS, targets might depend on other targets instead of cores depending on cores. Such an approach allows for fine-grained control of dependencies.
To declare target dependency, the user must call the hbs::AddDep procedure within the target procedure. The first argument is the dependency path. The remaining arguments are optional and are passed to the dependency procedure as arguments.
To add multiple distinct dependencies, the user must call hbs::AddDep multiple times. The ability to pass custom arguments to dependency was evaluated as much more advantageous than the ability to add multiple dependencies with a single hbs::AddDep call.
The hbs::AddDep internally calls the dependency procedure with the provided arguments. It also tracks dependencies so that generating a dependency graph is possible. Within a single flow, each target procedure can be run at most once with a particular set of arguments. This implies that if multiple target procedures add the same dependency with the same arguments, the dependency procedure is run only once during the first hbs::AddDep call. To enforce some target procedure rerun, the user can always directly call the target. However, enforcing target procedure rerun usually is an alert that a regular Tcl procedure shall be used instead of the core target procedure.
Listing 16 contains an example core definitions for presenting target dependency rules. There are four cores: core-a, core-b, core-c, generator-core. core-a depends on core-b and core-c. core-b depends on core-c. Moreover, cores core-a and core-b depends on the generator-core. However, they use different argument values for generation.
Listing 16: Example of target dependency rules.
namespace eval core-a {
proc target {} {
hbs::AddDep core-b::target
hbs::AddDep core-c::target
hbs::AddDep generator-core::gen a
hbs::AddDep generator-core::gen x
puts "core-a::target"
hbs::Register
namespace eval core-b {
proc target {} {
hbs::AddDep core-c::target
hbs::AddDep generator-core::gen b
hbs::AddDep generator-core::gen x
puts "core-b::target"
hbs::Register
namespace eval core-c {
proc target {} {
puts "core-c::target"
hbs::Register
namespace eval generator-core {
proc gen {arg} {
puts "generator-core::gen \$arg"
hbs::Register
Listing 17 presents output from running target core-a::target. As can be seen, the core-c::target is run only once, even though both core-a::target and core-b::target depend on it. This is because core-c::target has no arguments and can be added as a dependency only once. On the other hand, generator-core::gen is run three times. This is because generator-core::gen is added as a dependency four times, and three times with a distinct argument value.
Listing 17: Running target dependency example result.
[user@host tmp]\$ hbs run core-a::target
core-c::target
generator-core::gen b
generator-core::gen x
core-b::target
generator-core::gen a
core-a::target
4.9. EDA tool flow and stages
The primary function of any hardware build system is to provide the ability to build a design. What the term “build” actually means usually depends on the EDA tool. For some, it is only synthesis; for others, it is synthesis and implementation, and yet for others it might be simulation or just code linting. Each EDA tool is characterized by a distinct flow consisting of different stages. To mimic this behavior, HBS supports the following stages (alphabetical order):
  1. (1)
    analysis - HDL files analysis,
  2. (2)
    bitstream - device bitstream generation,
  3. (3)
    elaboration - design elaboration,
  4. (4)
    implementation - design implementation,
  5. (5)
    project - project creation,
  6. (6)
    simulation - design simulation,
  7. (7)
    synthesis - design synthesis.
Not all stages make sense for all EDA tools. Which stages are present in the given tool flow depends on that tool implementation. Namely, on the implementation of the hbs::<tool>::run procedure. This procedure is called at the end of the hbs::Run procedure execution.
Figure 1 shows the structure of the tool flow for the Vivado project mode. There are four stages: project, synthesis, implementation and bitstream. Each stage is wrapped by custom, user-defined callbacks. The number of callbacks is unlimited. Callbacks in any pre or post stage are executed in the order users add them. The user adds a callback by calling a dedicated HBS Application Programming Interface (API) procedure. For example, to add a post synthesis callback user can call the hbs::AddPostSynthCb procedure. Callbacks can be added with custom argument values.
The post stage callbacks and pre stage callbacks for adjacent stages were added for two reasons.
  1. (1)
    To clearly communicate which stage the callback refers to. For example, configuring implementation settings based on the synthesis results can be done in a post-synthesis callback or pre-implementation callback. However, as the callback modifies the implementation settings, it is probably better to add it using the hbs::AddPreImplCb, than hbs::AddPostSynthCb.
  2. (2)
    To introduce, to some extent, a manageable order of callbacks execution. The pre and post callbacks of a given stage are executed in the order they are added. If some nested dependencies add callbacks for the same pre or post stage, then the order of callbacks execution depends on the order of hbs::AddDep calls. However, if the result of one of the callbacks depends on the result of the other one, then relying on a user to call the hbs::AddDep procedures in proper order is error prone. In such a case, the callback that must be executed as the first one can be added to the post-synthesis callbacks, and the second callback can be added to the pre-implementation callbacks. Such an approach is immune to the order of hbs::AddDep calls.
The user can utilize stage callbacks in any desired way. However, the primary purpose of stage callbacks is to adjust the design build based on the results from a particular stage. For example, the user might want to configure additional implementation settings based on the synthesis results. The user might even terminate the tool flow in a given callback and report an error if certain conditions are not met.
Figure 1. Vivado project mode tool flow.
4.10. EDA tool commands custom arguments
The HBS has a minimal common abstraction layer. The user performs all the actions not covered by the abstraction layer by directly calling EDA tool commands. For example, generating extra timing or clock network reports. However, some of the actions are hidden under the HBS API. For example, adding or analyzing HDL files. EDA tool commands used to perform actions hidden under common API usually have some parameters that are not used by default. Nevertheless, sometimes there is a need to specify additional parameters. In such a case, there are two possible solutions.
The first option is to bypass the HBS API and directly call the underlying EDA tool command. The drawback of this approach is that the user must manually handle the current context. For example, when adding an HDL file, the user must manually specify the library or standard revision. Bypassing HBS API also bypasses the target context.
The second option is to set the underlying command arguments prefix or suffix. This can be achieved with the hbs::SetArgsPrefix and hbs::SetArgsSuffix procedures. The argument prefix is always appended after the command name, and the argument suffix is always appended after all command arguments.
Listing 18 presents Ethernet Management Data Input/Output (MDIO) core definition. The core has one testbench target utilizing the nvc simulator. Nvc report messages occupy multiple lines by default. However, this can be changed by specifying the --messages=compact argument when running the simulation. As running the simulation is the last stage of the nvc flow, the call to hbs::SetArgsPrefix must be wrapped by the call to the hbs::AddPostElabCb procedure.
Listing 18: Ethernet MDIO core definition.
namespace eval vhdl::ethernet {
namespace eval mdio {
proc src {} {
hbs::SetLib "ethernet"
hbs::AddFile mdio.vhd
proc tb {} {
hbs::SetTool "nvc"
hbs::AddPostElabCb hbs::SetArgsPrefix "--messages=compact"
src
hbs::SetLib ""
hbs::AddFile tb/tb-mdio.vhd
hbs::SetTop "tb_mdio"
hbs::Run
hbs::Register
Listing 19 shows commands executed by the hbs to run the simulation. The --messages=compact argument was appended right after the nvc command.
Listing 19: Ethernet MDIO testbench commands.
[user@host vhdl-ethernet]\$ hbs run vhdl::ethernet::mdio::tb
nvc --std=2019 -L. --work=ethernet -a vhdl/vhdl-ethernet/mdio.vhd
nvc --std=2019 -L. --work=work -a vhdl/vhdl-ethernet/tb/tb-mdio.vhd
nvc --std=2019 -L. -e tb_mdio
nvc --messages=compact --std=2019 -L. -r tb_mdio --wave
4.11. HBS API extra symbols
The HBS API consists not only of symbols related to the common EDA abstraction layer. For example, there are extra hbs::Exec and hbs::CoreDir procedures. The first one is a wrapper for the Tcl standard exec procedure. Before calling exec, the hbs::Exec changes the working directory to the directory where the currently evaluated core is defined. When exec returns, the hbs::Exec restores the working directory. The hbs::CoreDir procedure allows the user to get the path of the directory in which the currently evaluated core is defined.
HBS also provides users with the following extra variables:
  1. (1)
    hbs::ThisCorePath - the path of the core which target is currently being run,
  2. (2)
    hbs::ThisTargetPath - the path of the target which is currently being run,
  3. (3)
    hbs::ThisTargetName - the name of the target which is currently being run,
  4. (4)
    hbs::TopCorePath - the path of the top core beign run,
  5. (5)
    hbs::TopTargetPath - the path of the top target being run,
  6. (6)
    hbs::TopTarget - the name of the top target being run,
  7. (7)
    hbs::TopTargetArgs - the list with command line arguments passed to the top target.
4.12. Code generation
Code generation in the hardware design domain is omnipresent, as it significantly speeds up the implementation process. For example, hardware-software co-design system on chip projects usually have some tool automatically generating register files (Werner et al., 2020). Even in pure FPGA designs, it is common to generate descriptions at the register transfer level from some higher-level programming abstraction (Cong et al., 2011; Nane et al., 2016). That is why it is important for any hardware build system to provide as simple mechanism for code generation as possible.
Some existing hardware build systems, for example FuseSoc, do not allow the direct calling of an arbitrary external program in arbitrary places during the build process. Instead, the user has to define so-called generators (O., 2025). Only then can the user call the generator wihtin core definitions. However, such an approach has some drawbacks:
  1. (1)
    The generator call requires an extra layer of indirection. Generators are defined in different places than they are used, which decreases the readability of the description.
  2. (2)
    The generator call syntax does not resemble shell command call syntax. Generators are usually regular applications that can be executed in a shell. Calling a generator within the build system using syntax similar to shell seems natural.
In HBS, there is no formal concept of generator. Anything can be a generator, as generators are just regular Tcl procedures. This means that generators can be target procedures (tracked by dependency system) or core internal Tcl procedures (not tracked by the dependnecy system).
Listing 20 presents an example of calling an external code generator tracked by the dependnecy system. In actual usage, the call to the shell echo command would be replaced with a call to the proper code generator program. Calls to the hbs::AddFile are commented out because no EDA tool was set.
Listing 20: Example of HBS generator tracked by dependency system.
namespace eval core {
proc top {} {
hbs::AddDep generator::gen "foo"
puts "Adding file top.vhd"
# hbs::AddFile top.vhd
hbs::Register
namespace eval generator {
proc gen {name} {
exec echo "Generating \$name.vhd" >@ stdout
puts "Adding file \$name.vhd"
# hbs::AddFile "\$name.vhd"
hbs::Register
Listing 21 presents how to achieve the same result without tracking the generator as a dependency. This task is even more straightforward, as the user can call an external generator program directly in the target procedure.
Listing 21: Example of HBS generator not tracked by dependency system.
namespace eval core {
proc top {} {
exec echo "Generating foo.vhd" >@ stdout
puts "Adding file foo.vhd"
# hbs::AddFile "foo.vhd"
puts "Adding file top.vhd"
# hbs::AddFile top.vhd
hbs::Register
5. Hardware Build System Commands
The implemented hbs program provides the user with 11 commands. Listing 22 presents hbs help message. The help and version commands are not discussed in this section, as their purpose is rather obvious.
Listing 22: HBS help message.
Usage:
hbs <command> [arguments]
The command is one of:
doc Show documentation for hbs Tcl symbol or EDA tool
dump-cores Dump info about cores in JSON format
graph Output dependency graph for given target
help Print help message for program or given command
list-cores List cores
list-targets List targets
list-tb List testbench targets
run Run given target
test Run testbench targets
version Print hbs version
where Print where given cores are defined
Type hbs help <command>’ to obtain more
information about particular command.
5.1. Viewing HBS API documentation
The doc command was added to ease viewing documentatino for HBS Tcl symbols. Listing 23 presents an example of the doc command usage.
Listing 23: HBS doc command example.
[user@host tmp]\$ hbs doc AddFile
# AddFile adds files to the tool flow. Multiple files with different
# extensions can be added in a single call. args is the list of
# patterns used for globbing files. The file paths are relative to
# the ‘.hbs file path where the proc is called.
proc AddFile {args}
5.2. Dumping cores information
The dump-cores command allows dumping information about found cores into JSON format. The generated JSON data can be used for further processing. For example, the graph command utilizes data from JSON dump to generate the dependency graph.
5.3. Generating dependency graph
The graph command allows generating target dependency graphs in PDF format. Listing 24 shows the graph command help message.
Listing 24: HBS graph command help message.
Output dependency graph for given target.
hbs graph path/to/file/with/cores.json [top-target-path]
If top target path is not provided, then it is assumed to be the
same as the base of the json file path, without the ’.json suffix.
The graph command requires information about cores in the JSON format. This implies that the user must execute the dump-cores or run command before generating a dependency graph. However, this is not a major issue in practice. Dumping cores, even for large designs, does not take more than a few seconds, as the dump-cores command does not run the target tool flow.
Figure 2 presents an example dependency graph generated for VSC8211 (Ethernet PHY) chip tester design.
Figure 2. Example dependency graph generated by HBS graph command.
5.4. Listing cores
The list-cores command allows listing all cores discovered by the hbs program. Listing 25 shows list-cores command help message.
Listing 25: HBS list-cores command help message.
List cores found in .hbs files.
hbs list-cores [core-path-patterns...]
If core-path-patterns are not provided, all found cores are listed.
If core-path-patterns are provided, then only cores whose path
contains at least one pattern are listed.
For example, let us assume the user wants to list all cores in a project related to the APB bus. Listing 26 presents the output of such an action carried out in the VSC8211 tester design.
Listing 26: An example HBS list-cores command output.
[user@host tmp] hbs list-cores apb
vhdl::amba5::apb::bfm
vhdl::amba5::apb::checker
vhdl::amba5::apb::crossbar
vhdl::amba5::apb::mock-completer
vhdl::amba5::apb::pkg
vhdl::amba5::apb::serial-bridge
vhdl::amba5::apb::shared-bus
5.5. Listing targets
The list-targets command allows listing all targets discovered by the hbs program. Listing 27 shows the list-targets command help message. The list-targets command is analogous to the list-cores command but works on targets instead of cores.
Listing 27: HBS list-targets command help message.
List targets for cores found in .hbs files.
hbs list-targets [target-path-patterns...]
If target-path-patterns are not provided, all found targets
are listed. If target-path-patterns are provided, then only
targets whose path contains at least one pattern are listed.
For example, if you want to list all targets for a given core,
simply provide the core path as the argument.
For example, let us assume the user wants to list all targets of the APB serial bridge core. Listing 28 presents the output of such an action carried out in the VSC8211 tester design.
Listing 28: An example HBS list-targets command output.
[user@host tmp] hbs list-targets apb::serial
vhdl::amba5::apb::serial-bridge::src
vhdl::amba5::apb::serial-bridge::tb-write
vhdl::amba5::apb::serial-bridge::tb-read
5.6. Listing testbench targets
The list-tb command allows listing all testbench targets discovered by hbs. The list-tb command is analogous to the list-targets command, but it works on testbench targets instead of all targets. For example, let us assume the user wants to list all testbench targets of the APB serial bridge core. Listing 29 presents the output of such an action carried out in the VSC8211 tester design. As can be seen, the src target is not included.
Listing 29: An example HBS list-tb command output.
[user@host tmp] hbs list-tb apb::serial
vhdl::amba5::apb::serial-bridge::tb-write
vhdl::amba5::apb::serial-bridge::tb-read
5.7. Running target
The run command allows running target procedures. Usually, targets are run to carry out the build process or simulation. However, the user is free to carry out any action in the target being run. Listing 30 shows the run command help message. Running targets is described in subsections 4.5, 4.6, and 4.7.
Listing 30: HBS run command help message.
Run provided target.
hbs run <target-path> [target-args...]
The target path must be absolute target path containing the core
path and target name. Target arguments are forwarded to the target
procedure call.
Example:
hbs run core::path::target-name synthesis
5.8. Running testbench targets
The test command allows running all automatically discovered testbench targets. Listing 31 shows the test command help message.
Listing 31: HBS test command help message.
Run testbench targets.
hbs test [-workers N] [target-path-patterns...]
Testbench targets are detected automatically.
A testbench target is a target which name:
- starts with "tb-" or "tb_",
- ends with "-tb" or "_tb",
- equals "tb".
-workers specifies the number of targets to run simultaneously
(in parallel). N must be a positive integer. -workers must
be provided before target path patterns. Otherwise, it will
be treated as one of the target path patterns. By default,
the number of workers equals the number of available CPUs.
If target path patterns are not provided, all testbench targets
are run. If target path patterns are provided, then only testbench
targets whose path contains at least one pattern are run.
5.9. Locating core definition
The where command allows easily locating .hbs files in which given cores are defined. Listing 32 shows the where command help message.
Listing 32: HBS where command help message.
Print where given cores are defined.
hbs where [core-path-patterns...]
If core-path-patterns are not provided, all found cores are printed.
If core-path-patterns are provided, then only cores whose path
contains at least one pattern are printed.
For example, let us assume the user wants to locate the APB serial bridge core. Listing 33 presents the output of such an action carried out in the VSC8211 tester design.
Listing 33: An example HBS where command output.
[user@host tmp] hbs where apb::serial-bridge
vhdl::amba5::apb::serial-bridge /tmp/vsc8211-tester/gw/apb/apb.hbs
6. Hardware Build System Evaluation
This section tries to evaluate the implemented HBS in terms of other hardware build systems.
Comparing build systems is not straightforward. There are no objective numerical metrics that can be used for build systems comparison. One of such metrics could be the build time. However, in the case of hardware designs, the share of time consumed by EDA tools is orders of magnitude greater than the time overhead introduced by the build system code execution.
Some generic software evaluation metrics could be used, such as numerical or memory complexity. However, the memory overhead introduced by build systems is neglectable compared to the memory utilization of EDA tools. The numerical complexity might be important for the build system developers, but it is rather irrelevant for build system users.
This section tries to evaluate the HBS from some practical viewpoints, such as dependencies, overall internal complexity, ease of use, and provided functionalities.
6.1. Dependencies
The implemented HBS has only the following three dependencies:
  1. (1)
    Tcl shell (mandatory) - required to run the core logic of the build system.
  2. (2)
    Python (optional) - required to run commands not directly related to the build (automatic testbench running and dependency graph generation).
  3. (3)
    Graphviz (optional) - required for the graph command to generate a dependency graph in PDF format.
None of the existing hardware build systems capable of similar features has such a low number of dependencies.
6.2. Overall internal complexity
The measure used to estimate the overall internal complexity of the build system is the number of files, the number of lines of code, and the number of dependencies. The following type of files were excluded from analysis: author files, news files, contributors files, license files, readme files, tests files, and example files.
Table 1 presents metric values for the analyzed hardware build systems. The $\geq N$” signature means at least $N$. For build systems implemented in Python, the number of dependencies was analyzed using the pipreqs (V., 2025) tool. For build systems implemented in Rust, the number of dependencies was extracted from the Cargo.toml files. For flgen and bazel_hdl_rules, the number of dependencies was analyzed manually. Determining the exact number of dependencies (except for HBS) was not straightforward, as there might be additional runtime dependencies. Hence, the need for the greater-than-equal sign.
Besides the complexity metrics, the table also includes a summary of hardware build system features. The Introduced File Extensions is the number of custom file extensions used by a build system. The Introduced File Syntaxes is the number of custom syntaxes used by a build system. For example, FuseSoc scans for files with the .core extension. However, the .core files have a standard YAML syntax. This is why the number of introduced file extensions equals 1, but the number of introduced file syntaxes equals 0.
Table 1. Hardware build systems comparison (Y - yes, N - no, P - partially).
Build System HBS FuseSoc SiliconCompiler HAMMER Hog hdlmake bender orbit Blockwork flgen bazel_rules_hdl
Number of
Files
2 67 552 496 99 94 73 211 87 42 360
Lines of
Code
2639 6460 58543 136679 18982 8494 12026 46690 8610 3621 42743
External
Dependencies
3 >12 >37 >9 >3 >5 >23 >20 >16 >10 >7
Implementation
Language
Tcl Python Python Python Shell/Tcl Python Rust Rust Python Ruby Starlark
Introduced
File Extensions
1 1 0 0 4 1 2 1 0 1 0
Introduced
File Syntaxes
0 0 0 0 1 0 0 0 0 1 1
Dependency
Graph Generation
Y N Y N N N Y Y N N N
Build System
Code Executed
During
Design Build
Y N N N P N N N N N N
Build System
Code Executed
Directly by
EDA Tool
Y N N N P N N N N N N
Tied to Single
Revision System
N N N N Y N N N N N N
Requires Specific
Directory
Structure
N N N N Y N N N N N N
Capable of
Building Designs
Y Y Y Y Y Y Y N Y N Y
Requires
Containerization
N N N N N N N N Y N N
Automatic
Testbench
Running
Y N N N N N N N N N N
An extra context has to be introduced to the table, as different tools support different features and EDA tools.
Direct Tcl build systems (except HBS) are not included in the table. This is because they support only one EDA tool, and they do not support automatic testbench detection and dependency graph generation. OSVVM-Script is not included because it targets only simulation tools.
The FuseSoc build tool internally utilizes the external Edalize (Kindgren, 2025a) library for interaction with the EDA tools. Including Edalize, the number of files for FuseSoc equals 215, and the number of lines of code equals 19788.
The bender is only a dependency management tool. It provides a way to define dependencies among cores, execute unit tests, and verify that the source files are valid input for various simulation and synthesis tools. However, it does not run the EDA tool for building designs.
The flgen is just a file list generator. It does not run EDA tools for building designs.
Only HBS, SiliconCompiler, bender, and orbit are capable of generating dependency graphs.
Total number of lines of code of any hardware build system greatly depends on the number of supported EDA tools. FuseSoc supports the greatest number of tools, almost 40. HBS currently supports only five tools (AMD Vivado, AMD xsim, GHDL, GOWIN IDE, nvc). The most complex one is Vivado, which full support requires roughly 220 lines of code. However, supporting the nvc simulator required roughly 170 lines of code. Assuming that HBS will one day support 40 EDA tools, the number of lines of code should still be below 10000.
6.3. Ease of use
The subsection presents the ease of use of HBS in a very common task. Namely, the engineer implements a module and wants to scope some constraints to this module. As an example, a two-stage flip-flop clock domain crossing synchronizer is used. The source code of the synchronizer is not relevant to the analysis. False path is set as a constraint, although the max delay is also frequently used.
Listing 34 presents the synchronizer core definition in FuseSoc.
Listing 34: Cross domain crossing synchronizer core definition in FuseSoc.
CAPI=2:
name: vendor:library:synchronizer
filesets:
src:
files:
- synchronizer.vhd
file_type: vhdlSource-2008
vivado_constr:
files:
- vivado-constr.xdc: {file_type: xdc}
tcl:
files:
- constr.tcl: {file_type: tclSource}
targets:
default:
filesets:
- src
- "tool_vivado? (vivado_constr)"
- tcl
Listing 35 presents the content of the vivado-constr.xdc file and Listing 36 presents the content of the constr.tcl file.
Listing 35: Content of the vivado-constr.xdc file.
set_false_path -to [get_pins s_0_reg[*]/D]
Listing 36: Content of the constr.tcl file for FuseSoc.
if {[string match "Vivado*" [version]]} {
set_property SCOPED_TO_REF Synchronizer [get_files vivado-constr.xdc]
} else {
error "Synchronizer entity misses constraint file for your EDA tool"
In FuseSoc, to accomplish the task, the user must provide four files:
  1. (1)
    synchronizer.vhd - synchronizer HDL description,
  2. (2)
    vivado-constr.xdc - false path Vivado constraint for synchronizer,
  3. (3)
    constr.tcl - Tcl script scoping constraint to synchronizer module.
  4. (4)
    synchronizer.core - synchronizer core definition.
The number of required files is greater than the number of logical elements. The user describes the module in HDL, defines the core for the build system, and adds extra constraints. The number of logical elements equals three, not four. Moreover, the logic of the core definition is more complex to follow, as the user has to switch between files for a very short snippet of code. This leads to a decrease in readability.
Listing 37 presents the same core definition but in the HBS format. The number of required files equals three and is the same as the number of logical elements. The constr.tcl file could be eliminated because the code of the build system is executed directly by the EDA tool. Even the error message is improved. It is possible to report precisely for which EDA tool the constraint file definition is missing. Again, this is possible because the code of the build systems is executed directly by the EDA tool.
In the case of FuseSoc, build scripts are prepared before the EDA tool is run. Within the constr.tcl script, there is no access to the FuseSoc context and variables, as the build system code execution is already finished.
Listing 37: Cross domain crossing synchronizer core definition in HBS.
namespace eval vendor::library::synchronizer {
proc src {} {
hbs::AddFile synchronizer.vhd
if {\$hbs::Tool == "vivado-prj"} {
hbs::AddFile vivado-constr.xdc
set_property SCOPED_TO_REF Synchronizer [get_files vivado-constr.xdc]
} else {
error "Synchronizer entity misses constraint file for \$hbs::Tool"
hbs::Register
7. Conclusions
This work proposes a new Tcl-based hardware build system with the following features:
  1. (1)
    Low internal code complexity.
  2. (2)
    Minimal command line user interface.
  3. (3)
    Minimal Tcl user API.
  4. (4)
    Minimal common abstraction layer for EDA tools.
  5. (5)
    Minimal number of dependencies.
  6. (6)
    Ease of custom interaction with EDA tools.
  7. (7)
    Build system code executed directly by EDA tools during the build process.
  8. (8)
    Version control system agnosticity.
Although the proposed HBS represents a minimal approach in multiple aspects, it does not miss any significant features present in the existing build systems. It even supports functionalities absent in the existing systems.
While the implemented build system might not be the best choice for all kinds of projects, its unique set of features makes it appealing, especially for projects targeting multiple devices with multiple configurations.
The proposed Hardware Build System is novel because:
  1. (1)
    It is the only Tcl-based hardware build system supporting multiple EDA tools.
  2. (2)
    It is the only Tcl-based hardware build system capable of dependency graph generation, automatic testbench targets detection, and parallel testbench running.
  3. (3)
    It is the only hardware build system providing features described in bullets 1 and 2, allowing direct interaction with EDA tools during the build time. This is possible because the HBS code is directly executed by EDA tools.
Acknowledgements.
The work has been partially supported by the statutory funds of the Institute of Electronic Systems, Faculty of Electronics and Information Technology, Warsaw University of Technology.

### v4

@@
-HBS - Hardware Build System: A Tcl-based, minimal common abstraction approach for build system for hardware designs
+HBS - Hardware Build System: Characterizing and comparing direct-Tcl and indirect-abstract approaches for hardware build systems
 Abstract.
@@
 However, in the hardware description domain, no official build systems have been released with the predominant Hardware Description Languages (HDL) such as VHDL or SystemVerilog.
-Moreover, hardware design projects are often multilanguage.
-The paper proposes a new build system for the hardware description domain.
-The system is called the Hardware Build System (HBS).
-The main goals of the system include simplicity, readability, a minimal number of dependencies, and ease of integration with the existing Electronic Design Automation (EDA) tools.
-The system proposes a novel, minimal common abstraction approach, whose particular implications are described in the article.
-All the core functionalities are implemented in Tcl.
-Only the EDA tool’s independent features, such as dependency graph generation, are implemented in a Python wrapper.
+Moreover, hardware design projects are often multilingual.
+The paper characterizes and compares two common approaches for hardware build system implementations.
+The first one, the direct-Tcl approach, in which the build system code is executed directly by the EDA tool during the design build flow.
+The second one, the indirect-abstract approach, in which the build system produces a Tcl script, which is later run by a proper EDA tool.
+As none of the existing direct-Tcl build systems was close to the indirect-abstract build systems in terms of supported functionalities, the paper also presents a new direct-Tcl hardware build system called HBS.
+The implemented build system was used as a representative of direct-Tcl build systems in comparison with indirect-abstract build systems.
 Automation, Build System, Hardware Synthesis, Project Maintenance, Testbench Runner
@@
 Equally important is defining metadata used for configuration, dependency management, code generation, testing, and deployment purposes.
-The metadata is read by software tools, which, in turn, are responsible for generating final binary or bitstream files, or for carrying out extra tasks such as unit tests.
+The metadata is read by software tools, which, in turn, generate final binary or bitstream files or perform additional tasks, such as unit tests.
 Such programs are often referred to as build systems.
@@
 An introduction to the build system concept and principles can be found in (Smith, 2011; MOKHOV et al., 2020).
-In the software domain, there are multiple well-established build systems, such as GNU Make (Free Software Foundation, 2025), CMake (CMake, 2025), Ninja (Martin, 2025), or Bazel (Google, 2025).
+In the software domain, there are multiple well-established build systems, such as GNU Make (Free Software Foundation, 2025), CMake (CMake, 2025), Ninja (Martin, 2025), or Bazel (Mithro and others, 2021; Google, 2025).
 Some of them were implemented in the early days of programming.
 For example, GNU Make was first released in 1988 as part of the GNU Project.
-It was implemented to provide a free software alternative to Unix Make, which was first implemented in 1976 by Stuart Feldman at Bell Labs.
-Modern programming languages, such as Go, Rust, or Zig, come with build systems included in the language toolchains.
+It was implemented to provide a free software alternative to Unix Make, which was first implemented in 1976 by Stuart Feldman at Bell Labs (Feldman, 1979).
+Modern programming languages, such as Go, Rust, or Zig, include build systems in their toolchains.
 The Zig build system can even be used as a build system for C/C++ projects.
@@
 EDA vendors did not (and still do not) cooperate to create a standard (or at least similar) user interface for tool interaction.
-The main common point between various EDA tools is using the Tcl language as the interpreter for interaction with the tools.
+The main common point among various EDA tools is the use of the Tcl language as the interpreter for interacting with the tools.
 However, the tools provide custom commands and parameters even for primary, common tasks, such as adding hardware description files to a project design.
 For example, to add a VHDL file in AMD Vivado, one can use specific read_vhdl <path_to_vhdl_file> or generic add_files <path_to_vhdl_file> commands.
-To add a VHDL file in Intel Quartus, one has to use set_global_assignment -name VHDL_FILE <path_to_vhdl_file> command.
+To add a VHDL file in Intel/Altera Quartus, one has to use set_global_assignment -name VHDL_FILE <path_to_vhdl_file> command.
 EDA tools have abundant differences, making implementing a build system for hardware projects harder than for software.
-Despite this, there were (and still are) attempts to implement build systems that are more suitable for hardware projects.
-The paper proposes a new build system for a hardware description domain.
-The system is Tcl-based, which makes it easy to integrate with all existing EDA tools already utilizing Tcl language as a user-tool interface.
-A novel, minimal common abstraction approach differentiates the implemented system from the existing build systems targeting the hardware domain.
-The structure of the article is as follows.
-Section 2 describes related work.
-Section 3 discusses the requirements and their reasons imposed on the implemented build system.
-Section 4 introduces the internal architecture of the implemented build system.
-Section 5 describes commands provided by the implemented build system.
-Section 6 evaluates the proposed solution in terms of other build systems.
-Section 7 summarizes the work and highlights novel aspects of the implemented build system.
-2. Related work
-There already exist at least more than ten build systems targeting the hardware domain.
-Some of them are vendor tool specific, and some try to be generic, entirely abstracting away the tool used for synthesis or implementation.
+Despite this, there have been (and still are) attempts to implement build systems better suited to hardware projects.
+The main goal of the paper is to characterize and compare two existing approaches for hardware build system implementation, the direct-Tcl and indirect-abstract.
+The paper also partially describes the new hardware build system (HBS), which follows the direct-Tcl approach.
+The system was designed and implemented to serve as a comparison representative.
+The article’s structure is as follows.
+Section 2 describes two commonly used approaches for implementing hardware build systems.
+Section 3 explains why HBS was implemented.
+Section 4 describes requirements imposed on the HBS implementation and user interface.
+Section 5 briefly introduces the internal architecture of HBS.
+Section 6 compares various aspects of the direct-Tcl and indirect-abstract approaches.
+Section 7 summarizes the work.
+2. Hardware build system approaches
+There exist at least more than ten build systems targeting the hardware domain.
+Some of them are vendor-tool-specific, while others aim to be generic, entirely abstracting away the tool used for synthesis or implementation.
 By looking at the architecture (analyzing how a tool internally works), existing hardware build systems can be classified into two approaches:
   1. (1)
-    direct Tcl approach,
+    Direct-Tcl approach - the build system code is executed directly by the EDA tool during the design build flow.
+The build system code and the build code reside in the same domain.
+The build system code has access to the build code and vice versa.
   2. (2)
-    indirect abstract approach.
-2.1. Direct Tcl approach
-In the Tcl approach, the user directly utilizes the Tcl interface.
+    Indirect abstract approach - the build system code produces a Tcl script, which is only later run by a proper EDA tool.
+The build system code and the build code reside in different domains.
+The build system code generates the build code, but it does not have access to it during the design build flow.
+The build code is not aware of the build system code and has no access to it.
+2.1. Direct-Tcl approach
+In the direct-Tcl approach, the user directly writes Tcl code and uses the EDA tools’ Tcl interface.
 The idea seems obvious.
 Since EDA tools already utilize Tcl as a user interface, a build system should also provide users with a Tcl interface.
-Usually, the user does not directly call EDA tool specific commads.
-Tcl approch build systems implement some wrapper commands that the user should use.
-Representatives of the direct Tcl approach are, for example, fpga-vbs (Braun, 2025), vextproj (Zabołotny, 2025, 2016), and OSVVM-Scripts (Lewis, 2025).
-Being able to easily execute arbitrary, custom EDA tool commands is the most significant advantage of the direct Tcl approach.
-As a Tcl interpreter embedded into an EDA tool executes the build system code, a user can call custom commands in almost any place.
-The direct Tcl approach is not free of drawbacks. Some of them are inherent, and some result from the fact that direct Tcl build systems development was stopped at some stage by their maintainers.
+Usually, the direct-Tcl approach build systems implement some wrapper commands that the user should use for primary tasks, such as, file addition, library setting, or generic/parameter setting.
+However, in the case of more specific tasks, the user is supposed to call the EDA tool’s specific commands explicitly.
+It is also up to the user to maintain build compatibility across various EDA tools when build requirements are more complex, for example, when constraints depend on synthesis results.
+It is often achieved with tool guard clauses.
+For example, an if statement block calling a specific set of commands only if the currently executed EDA tool is specific.
+Representatives of the direct-Tcl approach are, for example, fpga-vbs (Braun, 2025), vextproj (Zabołotny, 2025, 2016), and OSVVM-Scripts (Lewis, 2025).
+Easily executing arbitrary, custom EDA tool commands is the most significant advantage of the direct-Tcl approach.
+As a Tcl interpreter embedded in an EDA tool executes the build system code, a user can call custom commands almost anywhere.
+The direct-Tcl approach is not free of drawbacks. Some of them are inherent, and others result from the fact that direct-Tcl build systems development was stopped at some stage by their maintainers.
 The first inherent drawback is that Tcl is considered an unfriendly language compared to, for example, Python.
@@
 The second inherent drawback is that Tcl is not a popular language nowadays.
-Although there are no physical obstacles to change this trend, it is doubtful that suddenly more developers start using Tcl instead, for example, Python as an interpreted language.
+Although there are no physical obstacles to change this trend, it is doubtful that suddenly more developers will start using Tcl instead of, for example, Python as an interpreted language.
 This is why this drawback is also treated as inherent.
-The second group of direct Tcl build systems drawbacks results from stopping the development at a certain stage.
-There is no direct Tcl build system that can be used to examine the limits and bottlenecks of this idea.
-For example, most of the direct Tcl build systems support only one EDA tool.
+The second group of direct-Tcl build systems drawbacks results from stopping development at a certain point.
+There is no direct-Tcl build system that can be used to examine the limits and bottlenecks of this idea.
+For example, most of the direct-Tcl build systems support only one EDA tool.
 OSVVM-Scripts is an exception.
@@
 It does not support any synthesis or implementation tools.
-More example drawbacks common to the direct Tcl build systems include:
+More example drawbacks common to the direct-Tcl build systems include:
   1. (1)
@@
   2. (2)
-    Lack or limited capability of automatic testbench targets detection.
+    Lack or limited capability of automatic cores, targets, and testbenches detection.
   3. (3)
     Lack of automatic parallel testbench running.
-2.2. Indirect abstract approach
-The indirect abstract approach is much more popular than the direct Tcl approach.
-In this approach, the user declares various properties and configurations instead of directly writing Tcl code or calling EDA tools Tcl commands.
-The most popular hardware build system utlizing the indirect abstract approach is probably FuseSoc (Kindgren, 2019, 2025b).
-Other popular build systems include SiliconCompiler (Olofsson et al., 2022; sil, 2025), HAMMER (Liew et al., 2022; ham, 2025), Hog (Biesuz et al., 2021; Gonnella et al., 2025), hdlmake (hdl, 2025), orbit (Ruskin, 2025), bender (ben, 2025), bazel_rules_hdl (baz, 2025), flgen (Ishitani, 2025), and Blockwork (blo, 2025).
-The language and syntax used for declarations vary between different build systems.
+The HBS was designed and implemented to examine the limits and bottlenecks of the direct-Tcl approach.
+To the author’s knowledge, there is no direct-Tcl approach hardware build system, supporting multiple EDA tools, with features and capabilities similar to those provided by the HBS.
+2.2. Indirect-abstract approach
+The indirect-abstract approach is much more popular than the direct-Tcl approach.
+In this approach, the user declares various properties and configurations rather than writing Tcl code or calling EDA tools’ Tcl commands directly.
+The main characteristic of the indirect-abstract approach is the execution of the build system code before the actual EDA tool is run.
+The result of the build system execution is a Tcl, Makefile, or a shell script, which is only later run by the specified EDA tool.
+The most popular hardware build system utilizing the indirect-abstract approach is probably FuseSoc (Kindgren, 2019, 2025c).
+Other popular build systems include SiliconCompiler (Olofsson et al., 2022; SiliconCompiler Project, 2025), HAMMER (Liew et al., 2022; Berkeley Architecture Research, 2025), Hog (Biesuz et al., 2021; Gonnella et al., 2025), hdlmake (CERN Open Hardware Repository, 2025), orbit (Ruskin, 2025), bender (PULP Platform, 2025), bazel_rules_hdl (The Hardware Description Languages GitHub Project, 2025), flgen (Ishitani, 2025), and Blockwork (Blockwork EDA, 2025).
+The indirect-abstract hardware build systems can be subdivided into the following three subcategories based on the language and syntax they utilize:
+  1. (1)
+    Declarative - the design build flow is described in a fully declarative way utilizing markup or data serialization languages, for example, YAML.
+An example of the declarative indirect-abstract build system is FuseSoc.
+  2. (2)
+    Programmable - the design build flow is described using a fully functional programming language, for example, Python.
+An example of the programmable indirect-abstract build system is SiliconCompiler.
+  3. (3)
+    Mixed - the design build flow is partially described using the declarative way and partially programmable. An example of the mixed indirect-abstract build system is Hog.
+The language and syntax used for describing build flow vary between different build systems.
 Some use simple data serialization formats such as YAML (for example, FuseSoc or bender) or TOML (for example, orbit).
-Some come with a domain-specific language (for example, flgen), or use Python (for example, HAMMER) and Makefiles (for example, Hdlmake).
+Some come with a domain-specific language (for example, flgen), or use Python (for example, SiliconCompiler, HAMMER) and Makefiles (for example, Hdlmake).
 Finally, some use a mix of more than one (for example, Hog).
-Especially using Python as a declarative language is powerful, as users can use all functionalities that come with the language.
-The most significant advantage of the indirect abstract approach is the low entry threshold.
-The syntax of languages such as YAML, TOML, or Python is simple, which leads to high legibility.
+Especially using Python as a programming language is powerful, as users can use all functionalities that come with the language.
+The most significant advantage of the indirect-abstract approach is the low entry threshold.
+The syntax of languages such as YAML, TOML, or Python is simple, which results in high legibility.
 Even junior engineers without extra training can understand it.
@@
 - src
-The indirect abstract approach is based on the concept that the interface to EDA tools is hidden under some common abstraction layer.
+The indirect-abstract approach is based on the concept that the interface to EDA tools is hidden under some common abstraction layer.
 Whether this is an advantage or a disadvantage is debatable.
 Providing a common abstract layer means that users do not have to learn the interface of multiple EDA tools.
-Moreover, in theory, changing the EDA tool for compilation process should be limited to changing solely the EDA tool declaration.
-However, this is somewhat a utopian view.
-As practice shows the common abstraction layer leads to the following common difficulties:
+Moreover, in theory, changing the EDA tool for the compilation process should be limited to changing the EDA tool declaration solely.
+However, this is somewhat of a utopian view.
+As the practice shows, the common abstraction layer leads to the following common difficulties:
   1. (1)
@@
 Let us suppose the abstraction layer declares feature F, and the build system claims support for tools T1 and T2.
-Although T1 might provide the implementation for F, T2 might be missing implementation for this particular feature.
+Although T1 might provide the implementation for F, T2 might be missing the implementation for this particular feature.
 In such a case, the user must add support for this feature to the build system.
 This task might not be trivial.
-The user must get to know the internals of the build system, prepare a pull request, wait for a review, and potentially reiterate the implementation.
+The user must understand the internals of the build system, prepare a pull request, wait for a review, and potentially reiterate the implementation.
 Yet another approach is to modify automatically generated build scripts before a compilation.
 However, this is against the whole idea of build systems.
-A direct Tcl approach is free of these difficulties because the user can directly insert custom Tcl commands.
+A direct-Tcl approach is free of these difficulties because the user can directly insert custom Tcl commands.
 There are no intermediate abstraction layers.
@@
 An example is discussed in subsection 6.4.
-All indirect abstract hardware build systems allow “injecting” arbitrary custom Tcl commands into the automatically generated scripts.
+All indirect-abstract hardware build systems allow “injecting” arbitrary custom Tcl commands into the automatically generated scripts.
 However, these commands can not be inserted into arbitrary places.
-Moreover, inserting a single-line Tcl command usually requires more than one line of code or even more than one file.
-This can be treated as a proof that the indirect abstract approach is not a golden bullet.
-The idea of a common abstraction layer has significant drawbacks that no one yet discovered how to fix.
-Making a truly abstract hardware build system is highlighy unliekly in the nearest feature.
+Moreover, inserting a single-line Tcl command often requires more than one line of code or even multiple files.
+This can be treated as proof that the indirect-abstract approach is not a golden bullet.
+The idea of a common abstraction layer has significant drawbacks that no one has yet found a way to fix.
+Making a truly abstract hardware build system is highly unlikely in the near future.
 The main problem with the common abstraction layer is the fact that EDA tools were implemented before such a layer was proposed.
-Declaring a common interface for already existing tools is simply much harder task than implementing tools conforming to some formal interface.
-The HBS described in the paper proposes a novel approach.
-It tries to incorporate the best of the direct Tcl approach and indirect abstract approach.
-It implements a common abstraction layer.
-However, the layer is:
+Declaring a common interface for existing tools is a much harder task than implementing tools that conform to a formal interface.
+3. Why Hardware Build System was implemented
+Characterizing and comparing direct-Tcl and indirect-abstract approaches for hardware build systems theoretically did not require HBS implementation.
+However, none of the existing direct-Tcl build systems was functionally close to the existing indirect-abstract build systems.
+In particular, none of the existing direct-Tcl build systems offered the following functionalities:
   1. (1)
-    minimal,
+    support for multiple synthesis EDA tools,
   2. (2)
-    limited to the primary common features,
+    listing automatically detected cores and targets,
   3. (3)
-    implemented in Tcl.
+    automatic testbench targets detection,
+  4. (4)
+    parallel testbench running,
+  5. (5)
+    dependency graph generation.
+The author wanted to implement a direct-Tcl build system that would functionally adhere to the existing indirect-abstract build systems.
+It had to be empirically verified if it is possible to have features of indirect-abstract build systems in a direct-Tcl build system.
+The implemented build system (HBS) serves as a representative of the direct-Tcl approach for the comparison with indirect-abstract build systems.
 4. Hardware Build System requirements
+The main goal of the HBS implementation was to push the direct-Tcl approach to its limits, examine its bottlenecks, and evaluate the concept of direct-Tcl build system by comparing it with indirect-abstract approaches.
+However, it does not mean that HBS is solely a prototype and was designed without thoughtful requirements.
+This section provides a summary of HBS major requirements with extra explanation.
 4.1. Simplicity and ease of use
@@
 The primary goal of the hardware build system proposed in the paper is simplicity.
-Contemporary hardware build systems utilizing the indirect abstract approach offer simplicity only at the initial stage.
-However, as the project complexity grows, so does the complexity of the build system utilization.
+Contemporary hardware build systems utilizing the indirect-abstract approach often offer simplicity only at the initial stage.
+However, as the project’s complexity grows, so does the complexity of using the build system.
 4.2. Minimal user interface
 Xia et al.
-in (Xia et al., 2013) report that even up to 21.35% of the bugs in software build systems belong to the external interface category.
-This means that one-fifth of build system bugs results from incorrect usage of the user interface or different behavior of the build system with different tools or platforms.
+(Xia et al., 2013) report that even up to 21.35% of the bugs in software build systems belong to the external interface category.
+This means that one-fifth of build system bugs result from incorrect usage of the user interface or different behavior of the build system with different tools or platforms.
 Fixing build system external interface bugs requires applying patches in the build system itself.
 Such a task is relatively time-consuming and might not be simple to implement.
-Build system user must report the problem to the build system developers or fix the bug himself.
-However, this implies that the build system user is distracted from his own work.
+The build system user must report the problem to the build system developers or fix the bug themselves.
+However, this implies that the build system user is distracted from their own work.
 Moreover, the fix is not applied immediately.
@@
 A minimal user interface has one more advantage.
-It allows developers to customize build systems to the project needs.
-Each project is in some way unique and might require particular approach for build problems solving.
-Moreover, projects usually evolve, for example, due to the customer demand.
+It allows developers to customize build systems to the project’s needs.
+Each project is in some way unique and might require a particular approach to solve problems related to the design build.
+Moreover, projects usually evolve, for example, due to customer demand.
 Dynamic tailoring to the project requirements is a desired feature of modern build systems (Maudoux and Mens, 2018).
@@
 Xiao et al.
-in (Xiao et al., 2022) report that dependency management is one of the top three factors of technical debt introduction in build systems.
+(Xiao et al., 2022) report that dependency management is one of the top three factors contributing to the introduction of technical debt in build systems.
 Such a finding is not surprising.
@@
 Getting a project to compile successfully is more important than fixing broken build system dependencies.
-One of the goals of the HBS proposed in the paper is to minimize the probability of user distraction caused by broken dependencies.
-This implies that the number of dependencies should be kept as low as possible.
+One of the goals of the HBS is to minimize the probability of user distraction caused by broken dependencies. This implies that the number of dependencies should be kept as low as possible.
 Moreover, a low number of dependencies potentially makes the build system more appealing for use in military designs.
@@
 For example, register files are automatically generated from SystemRDL (Accellera SYSTEMS INITIATIVE, 2025).
-In such a case, there must be a way to call an external program generating hardware description of the register components.
-Existing indirect abstract hardware build systems allow calling external programs.
+In such a case, there must be a way to call an external program that generates hardware description of the register components.
+Existing indirect-abstract hardware build systems allow calling external programs.
 However, they often require users to define a transitional so-called generator, which increases complexity and reduces legibility.
-Moreover, they usually do not allow to call external programs in arbitrary places.
+Moreover, they usually do not allow calling external programs in arbitrary places.
+This is further discussed in subsection 6.6.
 The second use case includes support for EDA tools without the Tcl interface.
@@
 The user simply calls programs with proper arguments and proper order in a shell.
-Being able to call arbitrary programs in arbitrary places is required to support this kind of tools without an extra script or Makefile generation.
-3.5. Independce from version control system
+Being able to call arbitrary programs in arbitrary places is required to support these kinds of tools without an extra script or Makefile generation.
+4.5. Independence from version control system
 Some hardware build systems are tightly coupled with a particular version control system.
@@
 5. Hardware Build System architecture
-This section describes the architecture of the implemented HBS. The system is available on GitHub (M., 2025) with the BSD 3-Clause License.
-It was decided that HBS will be implemented using Tcl language because of the following reasons:
+The HBS official user manual is available in its repository (M., 2025).
+The purpose of this section is to briefly describe the architecture of HBS so that the reader does not have to study the user manual.
+The HBS was implemented to evaluate the limitations of the direct-Tcl approach.
+Nevertheless, implementing a hardware build system using this approach may be a viable strategy for the following reasons:
   1. (1)
@@
 Moreover, these custom commands can be evaluated in arbitrary places.
-This aspect satisfies requirements 3.1 and 3.4.
-The complexity of existing indirect abstract build systems for less common tasks results from executing build system code before EDA tool flow.
+This aspect satisfies requirements 4.1 and 4.4.
+The complexity of existing indirect-abstract build systems for less common tasks results from executing build system code before the EDA tool flow.
 The build system only prepares scripts that are later run by EDA tools.
@@
 This implies that, in some cases, the build system user does not even have to install additional programs.
-This aspect satisfies requirement 3.3.
+This aspect satisfies requirement 4.3.
   3. (3)
-    Executing arbitrary programs in arbitrary places in Tcl is very simple.
+    Executing arbitrary programs in arbitrary places in Tcl is very straightforward.
 There is a dedicated exec command for invoking subprocesses.
 If executing a subprocess requires prior dynamic arguments evaluation, the exec command call must be prepended with the eval command.
-Even in Python, invoking a subprocess is not so straighforward.
-This Tcl feature satisfies requirement 3.4.
-One of the most important things while designing the HBS was the separation of common build process operations that would constitute a common abstraction layer over EDA tools.
-At the end, it was decided that the following actions should constitute the common abstraction layer:
-  1. (1)
-    Target device setting - some EDA tools use the term “part” instead of “device,” for example, Vivado.
-Simulation EDA tools do not require a device setting.
-However, all synthesis EDA tools require information on the target device.
-This is why setting the device became part of the common abstraction layer.
-  2. (2)
-    File addition - this includes support for adding files of all formats supported by a given EDA tool.
-  3. (3)
-    Library setting - setting HDL file library.
-  4. (4)
-    HDL standard setting - setting HDL file standard revision.
-The build system has to manage this because some tools can not handle analyzing different design units using different standard revisions, for example, nvc.
-In such a case, the build system must decide what the common standard revision shall be used for analyzing all HDL files.
-  5. (5)
-    Dependency specification - this is the core feature of any build system.
-  6. (6)
-    Generics/parameters setting - configuring parametric designs must be an inherent feature of any hardware build system.
-  7. (7)
-    Design top module setting - all EDA tools carrying out simulation or synthesis requires information on the top module.
-  8. (8)
-    Code generation - a hardware build system must provide a simple way for automatic arbitrary code generation. This is further explained in subsection 4.12.
-4.1. General structure
-The HBS is implemented in two files hbs.tcl and hbs.
-The first is implemented in Tcl, and the second in Python.
-The hbs.tcl file implements all the core features related directly to the interaction with the EDA tools.
-The hbs.tcl provides the following functions:
-  1. (1)
-    Dumping information about detected cores in JSON format.
-  2. (2)
-    Listing cores found in .hbs files.
-  3. (3)
-    Listing targets for a given core.
-  4. (4)
-    Running target provided as a command line argument.
-The hbs is a wrapper for the hbs.tcl and serves the following additional functions:
-  1. (1)
-    Showing documentation for hbs Tcl symbols.
-  2. (2)
-    Generating dependency graph.
-  3. (3)
-    Listing testbench targets.
-  4. (4)
-    Running testbench targets.
-By default, the user calls the hbs program.
-However, if none of the additional functions are required, the user can call the hbs.tcl directly.
-In such a case, the whole build system is limited to a single file.
-Check Table 1 to see how it compares to the size of existing hardware build systems.
+Even in Python, invoking a subprocess is not so simple.
+This Tcl feature satisfies requirement 4.4.
 5.1. Cores and cores detection
-When the user executes hbs (or hbs.tcl), all directories, starting from the working directory, are recursively scanned to discover all files with the .hbs extension (symbolic links are also scanned).
-Files with the .hbs extension are regular Tcl files that are sourced by the hbs.tcl script.
-However, before sourcing .hbs files, the file list is sorted so that scripts with shorter path depth are sourced as the first ones.
+When the user executes hbs, all directories under the working directory are recursively scanned to discover all .hbs files (including symbolic links).
+Files with the .hbs extension are regular Tcl files that are sourced by the hbs script.
+However, before sourcing .hbs files, the file list is sorted so that scripts with shorter path depth are sourced first.
 For example, let us assume the following three .hbs files were found:
@@
 Such an approach allows controlling when custom symbols (Tcl variables and procedures) are ready to use.
-For example, if the user has a custom procedure used in multiple .hbs files, then the user can create separate utils.hbs file contaning utility procedures, and place it in the the project root directory.
-Within .hbs files, the user usually defines cores and targets, although the user is free to have any valid Tcl code in .hbs files.
+For example, if the user has a custom procedure used in multiple .hbs files, then the user can create separate utils.hbs file containing utility procedures, and place it in the project root directory.
+Within .hbs files, the user usually defines cores and targets, though they are free to include any valid Tcl code.
 Listing 2 presents a very basic flip-flop core definition. The flip-flop core has a single target named src. The core consists of a single VHDL file.
@@
 hbs::Register
-To register a core, the user must explicitly call hbs::Register procedure at the end of the core namespace.
-Such a mechanism helps to distinguish regular Tcl namespaces from Tcl namespaces representing core definitions.
-If the user forgets to register a core, the build system gives a potential hint.
-An example is presented in Listing 3.
-Listing 3: An HBS error message with a hint on core registration.
-[user@host tmp]\$ hbs run lib::core::tb
-core lib::core not found, maybe the core is not registered (hsb::Register)
+To register a core, the user must explicitly call the hbs::Register procedure at the end of the core namespace.
+Such a mechanism helps distinguish regular Tcl namespaces from those representing core definitions.
 Each core is identified by its unique path.
@@
   1. (1)
-    The user can easily stick to the VLNV identifiers if required. This is presented in Listing 4. In this case, the flip-flop core path is vendor::library::flip-flop::1.0.
+    The user can easily stick to the VLNV identifiers if required. This is presented in Listing 3. In this case, the flip-flop core path is vendor::library::flip-flop::1.0.
   2. (2)
-    The user can define arbitrary deep core paths (limited by the Tcl shell). This is presented in Listing 5. In this case, the core path consists of 7 parts.
+    The user can define arbitrary deep core paths (limited by the Tcl shell). This is presented in Listing 4. In this case, the core path consists of 7 parts.
   3. (3)
     The user can nest namespaces to imitate the structure of libraries and packages.
-This is presented in Listing 6.
+This is presented in Listing 5.
 Three flip-flop cores are defined in the snippet.
-Listing 7 presents output for listing flip-flop cores.
-Listing 4: Example flip-flop core definition with VLNV core path.
+Listing 6 presents output for listing flip-flop cores.
+Listing 3: Example flip-flop core definition with VLNV core path.
 namespace eval vendor::library::flip-flop::1.0 {
@@
 hbs::Register
-Listing 5: Example flip-flop core definition with 7 core path parts.
+Listing 4: Example flip-flop core definition with 7 core path parts.
 namespace eval a::b::c::d::e::f::flip-flop {
@@
 hbs::Register
-Listing 6: An example .hbs file with three flip-flops core definitions
+Listing 5: An example .hbs file with three flip-flops core definitions
 namespace eval lib {
@@
 hbs::Register
-Listing 7: Listing flip-flop cores output.
-[user@host tmp]\$ hbs list-cores
+Listing 6: Listing flip-flop cores output.
+[user@host tmp]\$ hbs ls-cores
 lib::pkg1::d-flip-flop
@@
 5.2. Targets and targets detection
-HBS automatically detects targets.
-Targets are all Tcl procedures defined in the scope of core namespaces (namespaces with a call to hbs::Register).
+The HBS automatically detects targets.
+Targets are all Tcl procedures defined in the scope of core namespaces (namespaces with a call to the hbs::Register procedure).
 However, to allow users to define custom utility procedures within cores, procedures with names starting with the floor character (_) are not treated as core targets.
-Listing 8 presents an example edge detector core definition.
+Listing 7 presents an example edge detector core definition.
 The core path is vhdl::simple::edge-detector.
@@
 Such an approach allows for calling them in arbitrary places.
-The _tb procedure calls the src procedure because the edge_detector.vhd file is required for running testbench targets.
+The _tb procedure calls the src procedure because the edge_detector.vhd file is required for running the testbench targets.
 All targets are represented by a unique target path.
-Target path consists of the core path and target name.
+The target path consists of the core path and the target name.
 For example, the src target of the edge detector has the following path vhdl::simple::edge-detector::src.
-Listing 8: Example edge detector core definition.
+Listing 7: Example edge detector core definition.
 namespace eval vhdl::simple::edge-detector {
@@
 5.3. Testbench targets
-HBS is capable of automatically detecting testbench targets. Testbench targets are targets which names:
+The HBS is capable of automatically detecting testbench targets. Testbench targets are targets whose names:
   1. (1)
@@
     equal tb.
-For example, for the edge detector from Listing 8, the hbs program detects testbench targets presented in Listing 9.
-Listing 9: Testbench targets listed for example edge detector core.
-[user@host tmp]\$ hbs list-tb
+For example, for the edge detector from Listing 7, the hbs program detects testbench targets presented in Listing 8.
+Listing 8: Testbench targets listed for example edge detector core.
+[user@host tmp]\$ hbs ls-tb
 vhdl::simple::edge-detector::tb-comb
@@
 5.4. Running targets
-HBS allows running any target of registered cores.
-Even if the target itself has nothing to do with the hardware design.
-For example, running target print from Listing 10 results with the output shown in Listing 11.
-Listing 10: Example print target not related to hardware build.
+The HBS allows running any target of registered cores, even if the target itself has nothing to do with the hardware design.
+For example, running target print from Listing 9 results with the output shown in Listing 10.
+Listing 9: Example print target not related to hardware build.
 namespace eval core {
@@
 hbs::Register
-Listing 11: Running print target result.
+Listing 10: Running print target result.
 [user@host tmp]\$ hbs run core::print
@@
 However, in most cases, the user wants to run a target related to the flow of the set EDA tool.
-In such a case, instead of manually calling all of the required tool commands, the user can call the hbs::Run procedure in the core target procedure.
-The hbs::Run procedure has an optional argument accepting the stage after which the tool flow should stop.
-This is further described in subsection 4.9.
-After hbs::Run returns, the user can continue processing.
-For example, the user can run scripts analyzing code coverage or preparing additional reports.
+In such a case, instead of manually calling all required tool commands, the user can call the hbs::Run procedure from the core target procedure.
 5.5. Target parameters
@@
 Moreover, parameters can have optional default values.
-Additionally, HBS allows to provide command line arguments to the run target.
+Additionally, HBS allows providing command-line arguments to the run target.
 This is a very convenient feature in build systems.
-Listing 12 presents a very simplified example.
-Listing 12: Example target with an optional parameter.
+Listing 11 presents a very simplified example.
+Listing 11: Example target with an optional parameter.
 namespace eval core {
@@
 However, the example shows how the build stage can be passed from the command line to an EDA tool.
-Listing 13 presents output from running the target with different stage parameter values.
-Listing 13: Running print target result.
+Listing 12 presents output from running the target with different stage parameter values.
+Listing 12: Running print target result.
 [user@host tmp]\$ hbs run core::target
@@
 What target parameters are used for is limited only by the user’s imagination, and Tcl semantics.
-4.7. Target context
-An engineer implementing a given core controls the dependencies of the core.
-However, the core author does not control who will use the core and how.
-As targets are regular Tcl procedures, there is a need for a mechanism allowing the core author to evaluate the target procedure in the invariant environment.
-Such a mechanism in HBS is called the target context.
-The target context assures that the following variables are not affected by dependee or dependency target execution:
-  1. (1)
-    HDL library,
-  2. (2)
-    HDL standard,
-  3. (3)
-    top module name,
-  4. (4)
-    arguments prefix,
-  5. (5)
-    arguments suffix,
-  6. (6)
-    the core path,
-  7. (7)
-    the target name,
-  8. (8)
-    the target path.
-Listing 14 presents an example of the target context mechanism.
-Listing 14: Target context example.
-namespace eval pkg {
-namespace eval foo {
-proc src-foo {} {
-hbs::SetLib "lib-foo"
-hbs::AddDep pkg::bar::src-bar
-puts "foo lib: \$hbs::Lib"
-puts "foo core: \$hbs::ThisCorePath"
-puts "foo target: \$hbs::ThisTarget"
-hbs::Register
-namespace eval bar {
-proc src-bar {} {
-hbs::SetLib "lib-bar"
-puts "bar lib: \$hbs::Lib"
-puts "bar core: \$hbs::ThisCorePath"
-puts "bar target: \$hbs::ThisTarget"
-hbs::Register
-Listing 15 presents the output from running the pkg::foo:src-foo target. As can be seen, setting library in a target of one core, does not affect library in the target of another core.
-Listing 15: Output from target context example.
-[user@host tmp] hbs run pkg::foo::src-foo
-bar lib: lib-bar
-bar core: pkg::bar
-bar target: src-bar
-foo lib: lib-foo
-foo core: pkg::foo
-foo target: src-foo
-An author of a core can not easily by mistake break execution of another core target procedure. The target context provides the isolation of target procedure evaluation environments.
 5.6. Target dependencies
-In HBS, targets might depend on other targets instead of cores depending on cores. Such an approach allows for fine-grained control of dependencies.
+In HBS, targets depend on other targets rather than on cores. Such an approach allows for fine-grained control of dependencies.
 To declare target dependency, the user must call the hbs::AddDep procedure within the target procedure.
 The first argument is the dependency path.
-The remaining arguments are optional and are passed to the dependency procedure as arguments.
+The remaining arguments are optional and are forwarded to the dependency target procedure.
 To add multiple distinct dependencies, the user must call hbs::AddDep multiple times.
-The ability to pass custom arguments to dependency was evaluated as much more advantageous than the ability to add multiple dependencies with a single hbs::AddDep call.
+The ability to pass custom arguments to a dependency was evaluated as much more advantageous than the ability to add multiple dependencies with a single hbs::AddDep call.
 The hbs::AddDep internally calls the dependency procedure with the provided arguments.
-It also tracks dependencies so that generating a dependency graph is possible.
+It also tracks dependencies, enabling the generation of a dependency graph.
 Within a single flow, each target procedure can be run at most once with a particular set of arguments.
 This implies that if multiple target procedures add the same dependency with the same arguments, the dependency procedure is run only once during the first hbs::AddDep call.
-To enforce some target procedure rerun, the user can always directly call the target.
+To enforce a target procedure rerun, the user can always call the target procedure directly.
 However, enforcing target procedure rerun usually is an alert that a regular Tcl procedure shall be used instead of the core target procedure.
-Listing 16 contains an example core definitions for presenting target dependency rules.
-There are four cores: core-a, core-b, core-c, generator-core.
-core-a depends on core-b and core-c.
-core-b depends on core-c.
-Moreover, cores core-a and core-b depends on the generator-core.
-However, they use different argument values for generation.
-Listing 16: Example of target dependency rules.
-namespace eval core-a {
-proc target {} {
-hbs::AddDep core-b::target
-hbs::AddDep core-c::target
-hbs::AddDep generator-core::gen a
-hbs::AddDep generator-core::gen x
-puts "core-a::target"
-hbs::Register
-namespace eval core-b {
-proc target {} {
-hbs::AddDep core-c::target
-hbs::AddDep generator-core::gen b
-hbs::AddDep generator-core::gen x
-puts "core-b::target"
-hbs::Register
-namespace eval core-c {
-proc target {} {
-puts "core-c::target"
-hbs::Register
-namespace eval generator-core {
-proc gen {arg} {
-puts "generator-core::gen \$arg"
-hbs::Register
-Listing 17 presents output from running target core-a::target.
-As can be seen, the core-c::target is run only once, even though both core-a::target and core-b::target depend on it.
-This is because core-c::target has no arguments and can be added as a dependency only once.
-On the other hand, generator-core::gen is run three times.
-This is because generator-core::gen is added as a dependency four times, and three times with a distinct argument value.
-Listing 17: Running target dependency example result.
-[user@host tmp]\$ hbs run core-a::target
-core-c::target
-generator-core::gen b
-generator-core::gen x
-core-b::target
-generator-core::gen a
-core-a::target
-4.9. EDA tool flow and stages
-The primary function of any hardware build system is to provide the ability to build a design.
-What the term “build” actually means usually depends on the EDA tool.
-For some, it is only synthesis; for others, it is synthesis and implementation, and yet for others it might be simulation or just code linting.
-Each EDA tool is characterized by a distinct flow consisting of different stages.
-To mimic this behavior, HBS supports the following stages (alphabetical order):
-  1. (1)
-    analysis - HDL files analysis,
-  2. (2)
-    bitstream - device bitstream generation,
-  3. (3)
-    elaboration - design elaboration,
-  4. (4)
-    implementation - design implementation,
-  5. (5)
-    project - project creation,
-  6. (6)
-    simulation - design simulation,
-  7. (7)
-    synthesis - design synthesis.
-Not all stages make sense for all EDA tools.
-Which stages are present in the given tool flow depends on that tool implementation.
-Namely, on the implementation of the hbs::<tool>::run procedure.
-This procedure is called at the end of the hbs::Run procedure execution.
-Figure 1 shows the structure of the tool flow for the Vivado project mode.
-There are four stages: project, synthesis, implementation and bitstream.
-Each stage is wrapped by custom, user-defined callbacks.
-The number of callbacks is unlimited.
-Callbacks in any pre or post stage are executed in the order users add them.
-The user adds a callback by calling a dedicated HBS Application Programming Interface (API) procedure.
-For example, to add a post synthesis callback user can call the hbs::AddPostSynthCb procedure.
-Callbacks can be added with custom argument values.
-The post stage callbacks and pre stage callbacks for adjacent stages were added for two reasons.
-  1. (1)
-    To clearly communicate which stage the callback refers to.
-For example, configuring implementation settings based on the synthesis results can be done in a post-synthesis callback or pre-implementation callback.
-However, as the callback modifies the implementation settings, it is probably better to add it using the hbs::AddPreImplCb, than hbs::AddPostSynthCb.
-  2. (2)
-    To introduce, to some extent, a manageable order of callbacks execution.
-The pre and post callbacks of a given stage are executed in the order they are added.
-If some nested dependencies add callbacks for the same pre or post stage, then the order of callbacks execution depends on the order of hbs::AddDep calls.
-However, if the result of one of the callbacks depends on the result of the other one, then relying on a user to call the hbs::AddDep procedures in proper order is error prone.
-In such a case, the callback that must be executed as the first one can be added to the post-synthesis callbacks, and the second callback can be added to the pre-implementation callbacks.
-Such an approach is immune to the order of hbs::AddDep calls.
-The user can utilize stage callbacks in any desired way.
-However, the primary purpose of stage callbacks is to adjust the design build based on the results from a particular stage.
-For example, the user might want to configure additional implementation settings based on the synthesis results.
-The user might even terminate the tool flow in a given callback and report an error if certain conditions are not met.
-Figure 1. Vivado project mode tool flow.
-4.10. EDA tool commands custom arguments
-The HBS has a minimal common abstraction layer.
-The user performs all the actions not covered by the abstraction layer by directly calling EDA tool commands.
-For example, generating extra timing or clock network reports.
-However, some of the actions are hidden under the HBS API.
-For example, adding or analyzing HDL files.
-EDA tool commands used to perform actions hidden under common API usually have some parameters that are not used by default.
-Nevertheless, sometimes there is a need to specify additional parameters.
-In such a case, there are two possible solutions.
-The first option is to bypass the HBS API and directly call the underlying EDA tool command.
-The drawback of this approach is that the user must manually handle the current context.
-For example, when adding an HDL file, the user must manually specify the library or standard revision.
-Bypassing HBS API also bypasses the target context.
-The second option is to set the underlying command arguments prefix or suffix.
-This can be achieved with the hbs::SetArgsPrefix and hbs::SetArgsSuffix procedures.
-The argument prefix is always appended after the command name, and the argument suffix is always appended after all command arguments.
-Listing 18 presents Ethernet Management Data Input/Output (MDIO) core definition.
-The core has one testbench target utilizing the nvc simulator.
-Nvc report messages occupy multiple lines by default.
-However, this can be changed by specifying the --messages=compact argument when running the simulation.
-As running the simulation is the last stage of the nvc flow, the call to hbs::SetArgsPrefix must be wrapped by the call to the hbs::AddPostElabCb procedure.
-Listing 18: Ethernet MDIO core definition.
-namespace eval vhdl::ethernet {
-namespace eval mdio {
-proc src {} {
-hbs::SetLib "ethernet"
-hbs::AddFile mdio.vhd
-proc tb {} {
-hbs::SetTool "nvc"
-hbs::AddPostElabCb hbs::SetArgsPrefix "--messages=compact"
-src
-hbs::SetLib ""
-hbs::AddFile tb/tb-mdio.vhd
-hbs::SetTop "tb_mdio"
-hbs::Run
-hbs::Register
-Listing 19 shows commands executed by the hbs to run the simulation. The --messages=compact argument was appended right after the nvc command.
-Listing 19: Ethernet MDIO testbench commands.
-[user@host vhdl-ethernet]\$ hbs run vhdl::ethernet::mdio::tb
-nvc --std=2019 -L. --work=ethernet -a vhdl/vhdl-ethernet/mdio.vhd
-nvc --std=2019 -L. --work=work -a vhdl/vhdl-ethernet/tb/tb-mdio.vhd
-nvc --std=2019 -L. -e tb_mdio
-nvc --messages=compact --std=2019 -L. -r tb_mdio --wave
-4.11. HBS API extra symbols
-The HBS API consists not only of symbols related to the common EDA abstraction layer.
-For example, there are extra hbs::Exec and hbs::CoreDir procedures.
-The first one is a wrapper for the Tcl standard exec procedure.
-Before calling exec, the hbs::Exec changes the working directory to the directory where the currently evaluated core is defined.
-When exec returns, the hbs::Exec restores the working directory.
-The hbs::CoreDir procedure allows the user to get the path of the directory in which the currently evaluated core is defined.
-HBS also provides users with the following extra variables:
-  1. (1)
-    hbs::ThisCorePath - the path of the core which target is currently being run,
-  2. (2)
-    hbs::ThisTargetPath - the path of the target which is currently being run,
-  3. (3)
-    hbs::ThisTargetName - the name of the target which is currently being run,
-  4. (4)
-    hbs::TopCorePath - the path of the top core beign run,
-  5. (5)
-    hbs::TopTargetPath - the path of the top target being run,
-  6. (6)
-    hbs::TopTarget - the name of the top target being run,
-  7. (7)
-    hbs::TopTargetArgs - the list with command line arguments passed to the top target.
 5.7. Code generation
-Code generation in the hardware design domain is omnipresent, as it significantly speeds up the implementation process.
-For example, hardware-software co-design system on chip projects usually have some tool automatically generating register files (Werner et al., 2020).
-Even in pure FPGA designs, it is common to generate descriptions at the register transfer level from some higher-level programming abstraction (Cong et al., 2011; Nane et al., 2016).
-That is why it is important for any hardware build system to provide as simple mechanism for code generation as possible.
-Some existing hardware build systems, for example FuseSoc, do not allow the direct calling of an arbitrary external program in arbitrary places during the build process.
+Code generation in the hardware design domain is ubiquitous, as it significantly speeds up implementation.
+For example, hardware-software co-design system-on-chip projects usually have some tool automatically generating register files (Werner et al., 2020).
+Even in pure FPGA designs, it is common to generate register-transfer-level descriptions from a higher-level programming abstraction (Cong et al., 2011; Nane et al., 2016).
+That is why any hardware build system needs to provide as simple mechanism for code generation as possible.
+Some existing hardware build systems, for example, FuseSoc, do not allow the direct calling of an arbitrary external program ar arbitrary places during the build process.
 Instead, the user has to define so-called generators (O., 2025).
-Only then can the user call the generator wihtin core definitions.
+Only then can the user call the generator within core definitions.
 However, such an approach has some drawbacks:
   1. (1)
-    The generator call requires an extra layer of indirection. Generators are defined in different places than they are used, which decreases the readability of the description.
+    The generator call requires an extra layer of indirection. Generators are defined in different places than where they are used, which decreases the readability of the description.
   2. (2)
@@
 Calling a generator within the build system using syntax similar to shell seems natural.
-In HBS, there is no formal concept of generator.
+In HBS, there is no formal concept of a generator.
 Anything can be a generator, as generators are just regular Tcl procedures.
-This means that generators can be target procedures (tracked by dependency system) or core internal Tcl procedures (not tracked by the dependnecy system).
-Listing 20 presents an example of calling an external code generator tracked by the dependnecy system.
-In actual usage, the call to the shell echo command would be replaced with a call to the proper code generator program.
+This means that generators can be target procedures (tracked by the dependency system) or core internal Tcl procedures (not tracked by the dependency system).
+Listing 13 presents an example of calling an external code generator tracked by the dependency system.
+In actual use, the call to the shell’s echo command would be replaced with a call to the proper code-generation program.
 Calls to the hbs::AddFile are commented out because no EDA tool was set.
-Listing 20: Example of HBS generator tracked by dependency system.
+Listing 13: Example of HBS generator tracked by dependency system.
 namespace eval core {
@@
 hbs::Register
-Listing 21 presents how to achieve the same result without tracking the generator as a dependency.
+Listing 14 presents how to achieve the same result without tracking the generator as a dependency.
 This task is even more straightforward, as the user can call an external generator program directly in the target procedure.
-Listing 21: Example of HBS generator not tracked by dependency system.
+Listing 14: Example of HBS generator not tracked by dependency system.
 namespace eval core {
@@
 hbs::Register
-5. Hardware Build System Commands
-The implemented hbs program provides the user with 11 commands.
-Listing 22 presents hbs help message.
-The help and version commands are not discussed in this section, as their purpose is rather obvious.
-Listing 22: HBS help message.
-Usage:
+5.8. Hardware Build System Commands
+The implemented hbs program provides the user with 14 commands.
+Listing 15 presents hbs help message.
+Describing particular commands is considered out of scope of the article, as it would not bring any substantive value.
+The commands are listed solely to present the maturity and versatility of the implemented build system.
+None of the existing indirect-abstract approach hardware build systems support all features present in HBS.
+However, some have features missing in HBS, for example, SiliconCompiler, HAMMER, or Hog.
+Listing 15: HBS help message.
+Usage
 hbs <command> [arguments]
 The command is one of:
-doc Show documentation for hbs Tcl symbol or EDA tool
-dump-cores Dump info about cores in JSON format
+help Print help message
+doc Show documentation for cores
+dump Dump info about cores in Tcl dictionary format
+dump-json Dump info about cores in JSON format
 graph Output dependency graph for given target
-help Print help message for program or given command
-list-cores List cores
-list-targets List targets
-list-tb List testbench targets
+info Show information on hbs Tcl symbol or EDA tool
+ls-cores List cores found in .hbs files
+ls-targets List targets for given core
+ls-tb List testbench targets
 run Run given target
+dry-run Run given target without executing and evaluating commands
 test Run testbench targets
@@
 where Print where given cores are defined
-Type hbs help <command>’ to obtain more
-information about particular command.
-5.1. Viewing HBS API documentation
-The doc command was added to ease viewing documentatino for HBS Tcl symbols. Listing 23 presents an example of the doc command usage.
-Listing 23: HBS doc command example.
-[user@host tmp]\$ hbs doc AddFile
-# AddFile adds files to the tool flow. Multiple files with different
-# extensions can be added in a single call. args is the list of
-# patterns used for globbing files. The file paths are relative to
-# the ‘.hbs file path where the proc is called.
-proc AddFile {args}
-5.2. Dumping cores information
-The dump-cores command allows dumping information about found cores into JSON format.
-The generated JSON data can be used for further processing.
-For example, the graph command utilizes data from JSON dump to generate the dependency graph.
-5.3. Generating dependency graph
-The graph command allows generating target dependency graphs in PDF format. Listing 24 shows the graph command help message.
-Listing 24: HBS graph command help message.
-Output dependency graph for given target.
-hbs graph path/to/file/with/cores.json [top-target-path]
-If top target path is not provided, then it is assumed to be the
-same as the base of the json file path, without the ’.json suffix.
-The graph command requires information about cores in the JSON format.
-This implies that the user must execute the dump-cores or run command before generating a dependency graph.
-However, this is not a major issue in practice.
-Dumping cores, even for large designs, does not take more than a few seconds, as the dump-cores command does not run the target tool flow.
-Figure 2 presents an example dependency graph generated for VSC8211 (Ethernet PHY) chip tester design.
-Figure 2. Example dependency graph generated by HBS graph command.
-5.4. Listing cores
-The list-cores command allows listing all cores discovered by the hbs program. Listing 25 shows list-cores command help message.
-Listing 25: HBS list-cores command help message.
-List cores found in .hbs files.
-hbs list-cores [core-path-patterns...]
-If core-path-patterns are not provided, all found cores are listed.
-If core-path-patterns are provided, then only cores whose path
-contains at least one pattern are listed.
-For example, let us assume the user wants to list all cores in a project related to the APB bus. Listing 26 presents the output of such an action carried out in the VSC8211 tester design.
-Listing 26: An example HBS list-cores command output.
-[user@host tmp] hbs list-cores apb
-vhdl::amba5::apb::bfm
-vhdl::amba5::apb::checker
-vhdl::amba5::apb::crossbar
-vhdl::amba5::apb::mock-completer
-vhdl::amba5::apb::pkg
-vhdl::amba5::apb::serial-bridge
-vhdl::amba5::apb::shared-bus
-5.5. Listing targets
-The list-targets command allows listing all targets discovered by the hbs program.
-Listing 27 shows the list-targets command help message.
-The list-targets command is analogous to the list-cores command but works on targets instead of cores.
-Listing 27: HBS list-targets command help message.
-List targets for cores found in .hbs files.
-hbs list-targets [target-path-patterns...]
-If target-path-patterns are not provided, all found targets
-are listed. If target-path-patterns are provided, then only
-targets whose path contains at least one pattern are listed.
-For example, if you want to list all targets for a given core,
-simply provide the core path as the argument.
-For example, let us assume the user wants to list all targets of the APB serial bridge core. Listing 28 presents the output of such an action carried out in the VSC8211 tester design.
-Listing 28: An example HBS list-targets command output.
-[user@host tmp] hbs list-targets apb::serial
-vhdl::amba5::apb::serial-bridge::src
-vhdl::amba5::apb::serial-bridge::tb-write
-vhdl::amba5::apb::serial-bridge::tb-read
-5.6. Listing testbench targets
-The list-tb command allows listing all testbench targets discovered by hbs.
-The list-tb command is analogous to the list-targets command, but it works on testbench targets instead of all targets.
-For example, let us assume the user wants to list all testbench targets of the APB serial bridge core.
-Listing 29 presents the output of such an action carried out in the VSC8211 tester design.
-As can be seen, the src target is not included.
-Listing 29: An example HBS list-tb command output.
-[user@host tmp] hbs list-tb apb::serial
-vhdl::amba5::apb::serial-bridge::tb-write
-vhdl::amba5::apb::serial-bridge::tb-read
-5.7. Running target
-The run command allows running target procedures.
-Usually, targets are run to carry out the build process or simulation.
-However, the user is free to carry out any action in the target being run.
-Listing 30 shows the run command help message.
-Running targets is described in subsections 4.5, 4.6, and 4.7.
-Listing 30: HBS run command help message.
-Run provided target.
-hbs run <target-path> [target-args...]
-The target path must be absolute target path containing the core
-path and target name. Target arguments are forwarded to the target
-procedure call.
-Example:
-hbs run core::path::target-name synthesis
-5.8. Running testbench targets
-The test command allows running all automatically discovered testbench targets. Listing 31 shows the test command help message.
-Listing 31: HBS test command help message.
-Run testbench targets.
-hbs test [-workers N] [target-path-patterns...]
-Testbench targets are detected automatically.
-A testbench target is a target which name:
-- starts with "tb-" or "tb_",
-- ends with "-tb" or "_tb",
-- equals "tb".
--workers specifies the number of targets to run simultaneously
-(in parallel). N must be a positive integer. -workers must
-be provided before target path patterns. Otherwise, it will
-be treated as one of the target path patterns. By default,
-the number of workers equals the number of available CPUs.
-If target path patterns are not provided, all testbench targets
-are run. If target path patterns are provided, then only testbench
-targets whose path contains at least one pattern are run.
-5.9. Locating core definition
-The where command allows easily locating .hbs files in which given cores are defined. Listing 32 shows the where command help message.
-Listing 32: HBS where command help message.
-Print where given cores are defined.
-hbs where [core-path-patterns...]
-If core-path-patterns are not provided, all found cores are printed.
-If core-path-patterns are provided, then only cores whose path
-contains at least one pattern are printed.
-For example, let us assume the user wants to locate the APB serial bridge core. Listing 33 presents the output of such an action carried out in the VSC8211 tester design.
-Listing 33: An example HBS where command output.
-[user@host tmp] hbs where apb::serial-bridge
-vhdl::amba5::apb::serial-bridge /tmp/vsc8211-tester/gw/apb/apb.hbs
-6. Hardware Build System Evaluation
-This section tries to evaluate the implemented HBS in terms of other hardware build systems.
-Comparing build systems is not straightforward.
-There are no objective numerical metrics that can be used for build systems comparison.
-One of such metrics could be the build time.
+Type hbs help <command>’ to obtain more information about particular command.
+6. Direct-Tcl and indirect-abstract approaches comparison
+This section aims to characterize and compare direct-Tcl and indirect-abstract approaches for hardware build systems.
+The goal of the comparison is not to emerge the best hardware build system, but to present differences between build systems utilizing the direct-Tcl approach (build system code executed directly by EDA tools during the design build flow) and the ones utilizing the indirect-abstract approach (build system generates Tcl script which is later run by a proper EDA tool).
+The best choice of the hardware build system depends on the project or user requirements.
+Comparing hardware build systems is not straightforward.
+There are no objective numerical metrics that can be used for the build systems comparison.
+One such metric could be build time.
 However, in the case of hardware designs, the share of time consumed by EDA tools is orders of magnitude greater than the time overhead introduced by the build system code execution.
+This is why the build system execution time overhead is not taken into consideration at all.
 Some generic software evaluation metrics could be used, such as numerical or memory complexity.
-However, the memory overhead introduced by build systems is neglectable compared to the memory utilization of EDA tools.
+However, the memory overhead introduced by build systems is negligible compared to the memory utilization of EDA tools.
 The numerical complexity might be important for the build system developers, but it is rather irrelevant for build system users.
-This section tries to evaluate the HBS from some practical viewpoints, such as dependencies, overall internal complexity, ease of use, and provided functionalities.
-6.1. Dependencies
-The implemented HBS has only the following three dependencies:
-  1. (1)
-    Tcl shell (mandatory) - required to run the core logic of the build system.
-  2. (2)
-    Python (optional) - required to run commands not directly related to the build (automatic testbench running and dependency graph generation).
-  3. (3)
-    Graphviz (optional) - required for the graph command to generate a dependency graph in PDF format.
-None of the existing hardware build systems capable of similar features has such a low number of dependencies.
-6.2. Overall internal complexity
-The measure used to estimate the overall internal complexity of the build system is the number of files, the number of lines of code, and the number of dependencies.
-The following type of files were excluded from analysis: author files, news files, contributors files, license files, readme files, tests files, and example files.
-Table 1 presents metric values for the analyzed hardware build systems.
+This section provides a comparative analysis of direct-Tcl and indirect-abstract hardware build systems from a practical perspective.
+The qualitative assessment examines key operational attributes, including abstraction level, dependency overhead and its potential implications, internal code structure, constraints handling, handling EDA tools without Tcl interface, arbitrary external program calling, extra features, implementation complexity and learning difficulty.
+Table 1 presents some numerical and functional features for the analyzed hardware build systems.
 The $\geq N$” signature means at least $N$.
@@
 Hence, the need for the greater-than-equal sign.
-Besides the complexity metrics, the table also includes a summary of hardware build system features.
-The Introduced File Extensions is the number of custom file extensions used by a build system.
-The Introduced File Syntaxes is the number of custom syntaxes used by a build system.
-For example, FuseSoc scans for files with the .core extension.
-However, the .core files have a standard YAML syntax.
-This is why the number of introduced file extensions equals 1, but the number of introduced file syntaxes equals 0.
+The theory of operation of both approaches was analyzed in section 2, and is not included again in this section.
 Table 1. Hardware build systems comparison (Y - yes, N - no, P - partially).
@@
 Y N N N N N N N N N N
+Distributed
+execution of
+compilation flows
+N N Y N N N N N N N N
+6.1. Abstraction level
+The direct-Tcl and indirect-abstract approaches present entirely different paradigms for abstraction over EDA tool interfaces.
+In the direct-Tcl approach, the abstraction layer over EDA tool interfaces is limited to the common, primary actions.
+For example, in HBS, the following actions constitute the common abstraction layer over multiple EDA tools:
+  1. (1)
+    Target device setting. All synthesis EDA tools require information on the target device. This is why setting the device became part of the common abstraction layer.
+  2. (2)
+    File addition - this includes support for adding files of all formats supported by a given EDA tool.
+  3. (3)
+    Library setting - setting HDL file library.
+  4. (4)
+    HDL standard setting - setting HDL file standard revision.
+The build system has to manage this because some tools can not analyze different design units with different standard revisions, for example, nvc (Gasson, 2026).
+In such a case, the build system must decide which common standard revision to use for analyzing all HDL files.
+  5. (5)
+    Dependency specification - this is the core feature of any build system.
+  6. (6)
+    Generics/parameters setting - configuring parametric designs must be an inherent feature of any hardware build system.
+  7. (7)
+    Design top module setting - all EDA tools performing simulation or synthesis require information about the top module.
+  8. (8)
+    Exit severity setting - testbenches are designed to exit with an error code when a specific severity message occurs.
+This is independent of the simulator being used, hence it should be hidden under the build system abstraction.
+In the direct-Tcl approach, the build system intentionally exposes EDA tool-specific quirks so that the user has to resolve them explicitly.
+This is usually achieved with tool guard clauses.
+Listing 16 presents an example of such a guard.
+The snippet presents a helper testbench procedure definition for the AMBA 5 APB crossbar.
+If the tool is set to the nvc simulator, an extra argument ”–dump-arrays” must be passed to the simulator to dump array signals to the waveform file.
+This can be seen as a peculiarity of the nvc simulator.
+Listing 16: An example of tool guard clause in HBS (direct-Tcl approach).
+proc _tb {top} {
+hbs::AddDep vhdl::amba5::apb::checker::src
+hbs::AddDep vhdl::amba5::apb::bfm::src
+hbs::AddDep vhdl::amba5::apb::mock-completer::src
+if {\$hbs::Tool eq "nvc"} {
+hbs::AddPreSimCb hbs::SetArgSuffix "--dump-arrays"
+hbs::SetTop \$top
+hbs::AddDep \$hbs::ThisCorePath\::src
+In the indirect-abstract approach, the main purpose of the abstraction layer is to provide a great, unified abstraction over EDA tools.
+As such, the abstraction layer covers many more aspects than the abstraction layer in the direct-Tcl approach.
+However, maximizing the abstraction layer is not the goal of the indirect-abstract approach; the goal is to unify the user experience.
+Instead of exposing EDA tools peculiarities to the user, the indirect-abstract approach tries to hide them under the abstraction layer.
+For example, the nvc peculiarity of not dumping arrays signals by default, in the indirect-abstract approach is solved in such a way, that build systems automatically add the ”–dump-arrays” argument when simulation is run.
+6.2. Dependency overhead and its potential implications
+The term ”dependency” in the following context means a program, library, or operating system package that must be explicitly installed by the user to utilize all features of the analyzed build system.
+Any tools installed automatically alongside EDA tools are not treated as dependencies.
+Hardware build systems representing the direct-Tcl approach tend to have fewer dependencies. For example, the implemented HBS has only the following two dependencies:
+  1. (1)
+    Tcl shell (mandatory) - required to run the core logic of the build system.
+  2. (2)
+    Graphviz (optional) - required for generating a core dependency graph in a graphical format.
+None of the existing hardware build systems representing the indirect-abstract approach, capable of similar features, has so few dependencies.
+The explanation for the lower number of dependencies in build systems that use the direct-Tcl approach is straightforward.
+They rely on libraries and programs that are installed during EDA tool installation.
+On the other hand, build systems utilizing the indirect-abstract approach often rely on dependencies installed manually via package managers.
+Nevertheless, a pure number of dependencies might not be a good metric for comparison.
+A project could have a single dependency that suffers from multiple issues, and another project could have hundreds of well-maintained dependencies.
+What might be more important from the user’s point of view is stability.
+A build system is expected to work across many systems with a wide range of Long Term Support.
+The ease of initial setup of the build system might also be relevant to the user.
+The direct-Tcl approach with fewer dependencies is probably less susceptible to stability issues.
+However, overall stability depends on multiple factors, such as server stability, OS stability, network reliability, CI/CD complexity, etc.
+6.3. Internal code structure
+The internal code structure was analyzed in terms of the number of files, lines of code, introduced file syntaxes and extensions, and supported features.
+The following types of files were excluded from analysis: author files, news files, contributors files, license files, readme files, tests files, and example files.
+The goal of the analysis is to provide raw facts.
+In Table 1, the introduced file extensions are the number of custom file extensions used by a build system.
+The introduced file syntaxes are the number of custom syntaxes used by a build system.
+For example, FuseSoc scans for files with the .core extension.
+However, the .core files have a standard YAML syntax.
+This is why the number of introduced file extensions equals 1, but the number of introduced file syntaxes equals 0.
 An extra context has to be introduced to the table, as different tools support different features and EDA tools.
-Direct Tcl build systems (except HBS) are not included in the table.
+Direct-Tcl build systems (except HBS) are not included in the table.
 This is because they support only one EDA tool, and they do not support automatic testbench detection and dependency graph generation.
 OSVVM-Script is not included because it targets only simulation tools.
-The FuseSoc build tool internally utilizes the external Edalize (Kindgren, 2025a) library for interaction with the EDA tools.
+The FuseSoc build tool internally uses the external Edalize (Kindgren, 2025a) library to interact with EDA tools.
 Including Edalize, the number of files for FuseSoc equals 215, and the number of lines of code equals 19788.
-The bender is only a dependency management tool.
-It provides a way to define dependencies among cores, execute unit tests, and verify that the source files are valid input for various simulation and synthesis tools.
-However, it does not run the EDA tool for building designs.
-The flgen is just a file list generator. It does not run EDA tools for building designs.
-Only HBS, SiliconCompiler, bender, and orbit are capable of generating dependency graphs.
-Total number of lines of code of any hardware build system greatly depends on the number of supported EDA tools.
+The total number of lines of code of any hardware build system depends heavily on the number of supported EDA tools.
 FuseSoc supports the greatest number of tools, almost 40.
-HBS currently supports only five tools (AMD Vivado, AMD xsim, GHDL, GOWIN IDE, nvc).
-The most complex one is Vivado, which full support requires roughly 220 lines of code.
+HBS currently supports the following seven tools: AMD Vivado, AMD xsim, GHDL (Gingold, 2026), GOWIN IDE, Intel/Altera Quartus, nvc, Questa.
+The most complex one is Vivado, which full support requires roughly 190 lines of code.
 However, supporting the nvc simulator required roughly 170 lines of code.
-Assuming that HBS will one day support 40 EDA tools, the number of lines of code should still be below 10000.
-6.3. Ease of use
-The subsection presents the ease of use of HBS in a very common task.
-Namely, the engineer implements a module and wants to scope some constraints to this module.
+Assuming that HBS will one day support 40 EDA tools, the number of lines of code should remain below 10000.
+Tools following the direct-Tcl approach tend to have more monolithic code structure, fewer files, and fewer lines of code.
+The code for handling EDA tools is often placed in a single file.
+The lower number of lines of code results from the direct access to the EDA tool commands.
+There is no need for wrapper code.
+They prefer shared global variables which makes build logic easier to follow but also more bug-prone.
+Tools following the indirect-abstract approach tend to have a more modular code structure.
+They have more files and a greater number of lines of code.
+The code for handling specific EDA tools is organized into modules in separate files.
+The encapsulation makes the code more bug-resistant, but the build logic is slightly harder to follow.
+6.4. Constraints handling
+Building any hardware design (FPGA or ASIC) without constraints is impossible.
+Constrains usually define clocks, path delays, timing relations, and physical properties, for example, pin locations, drive strength or voltage level.
+Any hardware build system must support various constraint types and scenarios.
+The goal of this subsection is to present differences in constraint handling from the user perspective of the build system.
+In the case of a simple constraints scenario with global constraints, there is no difference between the direct-Tcl and indirect-abstract approaches.
+The user simply defines constraints in a supported file format (e.g., .sdc, .xdc, or .tcl) and adds the file to the project or build flow.
+Based on the file extension, build systems decide how to handle the file.
+However, there is a noticeable difference when a project requires constraints scoped to modules and cells, or when constraints are added optionally.
+Scoping constraints to module
+Let us analyze a simple example.
+The engineer implements a module and wants to scope some constraints to this module.
 As an example, a two-stage flip-flop clock domain crossing synchronizer is used.
 The source code of the synchronizer is not relevant to the analysis.
-False path is set as a constraint, although the max delay is also frequently used.
-Listing 34 presents the synchronizer core definition in FuseSoc.
-Listing 34: Cross domain crossing synchronizer core definition in FuseSoc.
+The false path is set as a constraint, although the max delay is also frequently used.
+Listing 17 presents the synchronizer core definition in FuseSoc, which represents the declarative indirect-abstract subcategory.
+Listing 17: Cross domain crossing synchronizer core definition in FuseSoc.
 CAPI=2:
@@
 - tcl
-Listing 35 presents the content of the vivado-constr.xdc file and Listing 36 presents the content of the constr.tcl file.
-Listing 35: Content of the vivado-constr.xdc file.
+Listing 18 presents the content of the vivado-constr.xdc file, and Listing 19 presents the content of the constr.tcl file.
+Listing 18: Content of the vivado-constr.xdc file.
 set_false_path -to [get_pins s_0_reg[*]/D]
-Listing 36: Content of the constr.tcl file for FuseSoc.
+Listing 19: Content of the constr.tcl file for FuseSoc.
 if {[string match "Vivado*" [version]]} {
@@
     synchronizer.core - synchronizer core definition.
-The number of required files is greater than the number of logical elements.
+The number of required files (4) is greater than the number of logical elements (3).
 The user describes the module in HDL, defines the core for the build system, and adds extra constraints.
-The number of logical elements equals three, not four.
-Moreover, the logic of the core definition is more complex to follow, as the user has to switch between files for a very short snippet of code.
-This leads to a decrease in readability.
-Listing 37 presents the same core definition but in the HBS format.
-The number of required files equals three and is the same as the number of logical elements.
+The number of logical elements equals three, and the number of files equals four.
+The logic of the core definition is more complex to follow, as the user has to switch between files for a very short snippet of code.
+This may lead to a decrease in productivity (Meyer et al., 2014).
+Listing 20 presents the same core definition but in the direct-Tcl approach.
+The number of required files is three and equals the number of logical elements.
 The constr.tcl file could be eliminated because the code of the build system is executed directly by the EDA tool.
 Even the error message is improved.
-It is possible to report precisely for which EDA tool the constraint file definition is missing.
-Again, this is possible because the code of the build systems is executed directly by the EDA tool.
-In the case of FuseSoc, build scripts are prepared before the EDA tool is run.
-Within the constr.tcl script, there is no access to the FuseSoc context and variables, as the build system code execution is already finished.
-Listing 37: Cross domain crossing synchronizer core definition in HBS.
+It is possible to report which EDA tool the constraint file definition is missing for.
+Again, this is possible because the build system code is executed directly by the EDA tool.
+In the case of the indirect-abstract approach, build scripts are prepared before the EDA tool is run.
+Within the constr.tcl script, there is no access to the build system context or variables, as the build system code execution is already finished.
+Listing 20: Cross domain crossing synchronizer core definition in HBS (direct-Tcl approach).
 namespace eval vendor::library::synchronizer {
@@
 hbs::Register
-7. Conclusions
-This work proposes a new Tcl-based hardware build system with the following features:
+In the case of the programmable indirect-abstract subcategory, the solution depends on the tool.
+For example, in hdlmake, the intermediate Tcl script is also required to scope a constraint file to a module.
+The same script is also required for the SiliconCompier.
+Hog, for example, instead of a Tcl script, requires a custom constraint list file (a file with the .con extension) with paths to constraint files and their property declarations.
+Moreover, the only way to opt out of the default constraints is to comment out the constraint lines in the constraint list file.
+To the author’s knowledge, none of the existing indirect-abstract build systems utilizes the concept of direct injection of Tcl code into the build script.
+They all need one additional file (a Tcl script or a tool-specific file) for scoping constraints.
+Scoping constraints to cells and optional constraints
+Let us analyze a more complex example.
+Namely, the Advanced Peripheral Bus Clock Domain Crossing bridge (APB CDC bridge).
+Such a bridge requires some timing constraints due to the clock domain crossing.
+There are a few possible ways to define valid constraints for the APB CDC bridge.
+The core designer may want to provide some default constraints.
+However, the user should also be able to provide their own constraints.
+One possible constraint strategy is to set a false path for handshake signals and a maximum delay for bus signals crossing domains.
+The max delay value should equal the period of the destination clock.
+The false path constraints can be safely scoped to the APB CDC bridge module.
+However, because there may be multiple APB CDC bridges in the design crossing between various clock domains, the max delay constraints should be scoped to cells.
+An alternative solution is to pick the shortest clock period as the value for the max delay constraint for all bridges.
+However, such an approach leads to design overconstraint and imposes unnecessary requirements on the router.
+Listing 21 shows how APB CDC bridge constraints are handled in the HBS, which represents the direct-Tcl approach.
+The code is concise and located within a single procedure.
+The user can pass the -no-constr flag to the target to omit adding the default constraints.
+The flag is checked with a simple if statement.
+Moreover, if an invalid argument is passed, an error with a precise message is returned.
+The apb-cdc-bridge-ref.xdc and apb-cdc-bridge-cell.tcl constraint files are not presented as their content is irrelevant.
+The first one has constraints scoped to the module; the second one has code that applies constraints scoped to cells.
+Listing 21: Advanced Peripheral Bus Clock Domain Crossing bridge target definition in HBS (direct-Tcl).
+# Use -no-constr flag to omit adding the default constraints.
+proc src {args} {
+hbs::AddDep vhdl::amba5::apb::pkg::src
+hbs::SetLib "lapb"
+hbs::AddFile cdc-bridge.vhd
+set argCount [llength \$args]
+if {\$argCount == 1} {
+set arg [lindex \$args 0]
+if {\$arg eq "-no-constr"} {
+return
+} else {
+error "\$hbs::ThisTargetPath: invalid arg ’\$arg’, valid args are: ’-no-constr’"
+} elseif {\$argCount > 1} {
+error "\$hbs::ThisTargetPath: invalid arg count \$argCount, expected at most 1"
+if {\$hbs::Tool eq "vivado-prj"} {
+hbs::AddFile vivado/apb-cdc-bridge-ref.xdc
+set_property SCOPED_TO_REF APB_CDC_Bridge [get_files apb-cdc-bridge-ref.xdc]
+hbs::AddFile vivado/apb-cdc-bridge-cell.tcl
+set_property USED_IN_SYNTHESIS false [get_files apb-cdc-bridge-cell.tcl]
+} else {
+error "\$hbs::ThisCore misses constraints for \$hbs::Tool"
+How applying the same constraints in the indirect-abstract approach looks depends on the tool subcategory.
+For example, in FuseSoc, which represents the declarative subcategory, applying the module and cell-scoped constraints would be similar to what was presented in listings 17 and 19 for the CDC synchronizer.
+Again, one more Tcl file would be required.
+The optionality of the constraints would be handled using the flags mechanism (Kindgren, 2025b).
+While the flag mechanism allows conditional inclusion of files, it does not allow arbitrary actions to be executed due to its declarative nature.
+In hdlmake, the same core description would require two Manifest.py files.
+The first one with constraints included, and the second one without.
+The user would pick a proper Manifest.py file when declaring dependency on the core.
+With just a single boolean parameter, this approach does not seem complex.
+However, with multiple parameters, the number of Manifest.py files would equal the number of possible combinations.
+In the programmable subcategory of the indirect-abstract approach, the complexity of constraint handling is lower.
+In this subcategory, the solution is very similar to the one present in the direct-Tcl approach.
+However, instead of directly injecting the Tcl code into the generated Tcl build script, the user must place it in a separate Tcl file.
+There is only minor overhead resulting from:
   1. (1)
-    Low internal code complexity.
+    build system programming language Application Programming Interface (API) call,
   2. (2)
-    Minimal command line user interface.
+    build logic fragmentation due to the requirement of putting the Tcl code in a separate file.
+6.5. Handling EDA tools without Tcl interface
+Not all EDA tools have a Tcl interface.
+For example, free and open-source simulators, for example, Verilator (Snyder, 2026) or GHDL, have a shell interface.
+The user simply calls programs within the shell providing appropriate arguments.
+Handling EDA tools without the Tcl interface in the direct-Tcl approach is straightforward.
+The Tcl language makes it easy to call any external program using the exec command.
+Listing 22 presents how the elaboration stage for the GHDL simulator is implemented in HBS.
+The command to be executed is constructed directly as a string to be passed to the shell.
+The code related to the working directory change and error checking takes more lines than the code creating the actual command.
+The hbs::Exec procedure is a simple wrapper around the Tcl exec command, which allows printing the command to the standard output in the dry-run mode.
+Listing 22: GHDL elaboration stage implementation in HBS (direct-Tcl approach).
+proc elaborate {} {
+set workDir [pwd]
+cd \$hbs::RunTargetBuildDir
+set cmd "ghdl -e \$hbs::ArgPrefix --std=[hbs::ghdl::std] --workdir=work \
+\$hbs::ghdl::libs \$hbs::ArgSuffix \$hbs::Top"
+set err [hbs::Exec \$cmd]
+if {\$err} {
+hbs::panic "\$hbs::Top elaboration failed with exit status \$err"
+cd \$workDir
+In the indirect-abstract approach, handling EDA tools without the Tcl interface is implemented differently.
+In the indirect-abstract approach, the build system code is run before the build code.
+This implies that the build system code must generate the build code that is only later run.
+The build code is usually generated in the form of a Makefile or shell script.
+Compared to the direct-Tcl approach, there are two essential differences:
   1. (1)
-    Minimal Tcl user API.
+    The generated build code script can be easily reused without rerunning the build system code.
   2. (2)
-    Minimal common abstraction layer for EDA tools.
+    The build system code logic might be slightly harder to follow.
+This is because the data for the script templates is usually not prepared in a linear way that corresponds to the build flow.
+Whereas in the direct-Tcl approach, the logic must be linear because the build code is executed during the build system code execution.
+6.6. Arbitrary external program calling
+Arbitrary external program calling is a very useful feature of modern hardware build systems. The word ”arbitrary” means two things:
   1. (1)
-    Minimal number of dependencies.
+    arbitrary program - the user should be able to call any program available on the operating system,
   2. (2)
-    Ease of custom interaction with EDA tools.
+    arbitrary place - the user should be able to call programs in any place within the design tool flow.
+A practical example of an arbitrary program is a code generator. Code generation in the hardware design domain is ubiquitous, as it significantly speeds up implementation.
+A practical example of an arbitrary place is report analysis.
+A user may want to analyze the synthesis stage results before starting the implementation stage.
+However, the user may want to implement the reports parsing logic in a Perl or Python script.
+In such a case, there must be a way to call the custom script within the Tcl build script after the synthesis.
+Meeting the two arbitrariness requirements (program and place) in the direct-Tcl approach is straightforward. This is because of the following reasons:
   1. (1)
-    Build system code executed directly by EDA tools during the build process.
+    Tcl has built-in exec command for executing external programs,
   2. (2)
-    Version control system agnosticity.
-Although the proposed HBS represents a minimal approach in multiple aspects, it does not miss any significant features present in the existing build systems.
-It even supports functionalities absent in the existing systems.
-While the implemented build system might not be the best choice for all kinds of projects, its unique set of features makes it appealing, especially for projects targeting multiple devices with multiple configurations.
-The proposed Hardware Build System is novel because:
+    the build system code is executed directly by the EDA tool,
   3. (3)
-    It is the only Tcl-based hardware build system supporting multiple EDA tools.
+    the build system code is executed during the design build flow.
+Listing 23 presents an example of how to execute a command after the synthesis step in HBS.
+Listing 23: Post synthesis callback example in HBS (direct-Tcl).
+proc myPostSynthCb {} {
+set err [catch {eval exec "./analyze-synth-reports.py"} output]
+if {\$output ne ""} {
+puts "synthesis results are not satisfactory: \$output"
+exit 1
+hbs::AddPostSynthCb myPostSynthCb
+In HBS, each stage has optional pre- and post-callbacks. Figure 1 shows the structure of the tool flow for the Vivado project mode.
+Figure 1. Vivado project mode tool flow.
+There are four stages: project, synthesis, implementation, and bitstream.
+The number of callbacks is unlimited.
+Callbacks in any pre- or post-stage are executed in the order users add them.
+The user adds a callback by calling a dedicated HBS API procedure.
+For example, to add a post-synthesis callback, the user can call the hbs::AddPostSynthCb procedure.
+Callbacks can be also added with custom argument values.
+While most existing indirect-abstract hardware build systems allow calling any third-party programs, the tools representing the declarative subcategory, do not allow calling them in arbitrary places.
+For example, in FuseSoc, the user has to define so-called generators (O., 2025).
+Only then can the user call the generator within core definitions.
+However, such an approach has some drawbacks:
   1. (1)
-    It is the only Tcl-based hardware build system capable of dependency graph generation, automatic testbench targets detection, and parallel testbench running.
+    The generator call requires an extra layer of indirection. Generators are defined in different places from where they are used, which reduces the readability of the description.
   2. (2)
-    It is the only hardware build system providing features described in bullets 1 and 2, allowing direct interaction with EDA tools during the build time.
-This is possible because the HBS code is directly executed by EDA tools.
+    The generator call syntax does not resemble shell command call syntax.
+Generators are usually regular applications that can be executed in a shell.
+Calling a generator within the build system using syntax similar to shell seems natural.
+  3. (3)
+    Generators can be executed only before the EDA tool flow.
+In the case of build systems representing the programmable indirect-abstract subcategory, meeting the two arbitrariness requirements is also straightforward. However:
+  1. (1)
+    it requires some Python wrapper code,
+  2. (2)
+    the code must be placed in a separate script file.
+Listing 24 shows how to add a post-synthesis script in SiliconCompiler.
+There are dedicated Python API functions, so the procedure is very concise.
+Compared to the direct-Tcl approach, there is one more indirection layer, as the actual Tcl code must be placed in a separate file, which leads to the build logic fragmentation.
+Listing 24: Post synthesis callback example in SiliconCompiler (indirect-abstract).
+SynthesisTask.find_task(project).add_postscript("path/to/post-synth-script.tcl")
+Listing 25 presents how to execute a command after the synthesis stage in hdlmake.
+The snippet is a part of a Manifest.py file.
+The hdlmake, like SiliconCompiler, requires the code to be placed in a separate Tcl file.
+Moreover, hdlmake closes the project after every step.
+If the evaluated code requires access to the build context, for example, synthesis results, the user must explicitly open the project in the script.
+Listing 25: Post synthesis program execution example in hdlmake (indirect-abstract).
+syn_post_synthesize_cmd = "\$(TCL_INTERPRETER) user_tcl_script.tcl;"
+6.7. Extra features support
+Modern build systems are expected to do much more than produce the final production or programming files. In terms of hardware build systems, for example, the following extra features are desired:
+  1. (1)
+    automatic testbench targets detection,
+  2. (2)
+    parallel testbench running,
+  3. (3)
+    core dependency graph generation,
+  4. (4)
+    distributed development and distributed execution of compilation flows.
+None of the existing direct-Tcl hardware build systems supports feature number 4. Moreover, HBS is the only direct-Tcl build system supporting features 1, 2, and 3.
+In the indirect-abstract class, only SiliconCompiler, bender, and orbit support feature number 3.
+The SiliconCompiler is the only existing hardware build system with built-in support for distributed development and distributed execution of compilation flows.
+SiliconCompiler and HAMMER have the most extensive support for ASIC and VLSI designs.
+Both were designed from the ground up for that purpose and used for commercial projects in this domain.
+However, both of them have the greatest internal complexity.
+It is also important to note that:
+  • The bender is only a dependency management tool.
+It provides a way to define dependencies among cores, execute unit tests, and verify that the source files are valid input for various simulation and synthesis tools.
+However, it does not run the EDA tool for building designs.
+  • The flgen is just a file list generator. It does not run EDA tools for building designs.
+6.8. Extra features implementation complexity
+While hardware build system users are mostly interested in extra-feature support (subsection 6.7), hardware build system developers might also be interested in the implementation complexity of extra features.
+Implementing extra, non-build-related features in the direct-Tcl approach is more complex than in the indirect-abstract approach.
+This state results directly from the use of the Tcl language.
+While multi-threading in Tcl is possible, the Tcl was not designed with multi-threading as one of its primary goals.
+This explicitly impacts support for parallel testbench execution or distributed compilation.
+Moreover, Tcl has fewer third-party libraries that could help implement additional features.
+Yet another important factor is the fact that direct-Tcl build systems are executed by the Tcl interpreters provided by the EDA tools.
+Each version of the EDA tool is shipped with a fixed version of the Tcl interpreter.
+Modern third-party Tcl libraries might require a higher version of the Tcl interpreter, which might potentially block them from being used within EDA tools.
+In the indirect-abstract approach, the build system infrastructure (tools and libraries) is managed independently from the EDA tools.
+The direct-Tcl approach tools tend to have more code related to the extra features implementation because any feature requiring parallelism has to be implemented asynchronously using the filevent and vwait mechanisms.
+Moreover, the code has a non-linear structure and is spread across multiple locations.
+An alternative approach to mitigate the mentioned deficiencies in the direct-Tcl approach is to implement the extra features in a wrapper program or script written in another programming language.
+However, this leads to splitting the build system’s code logic across multiple languages, which might increase cognitive complexity.
+Hardware build systems that follow the indirect-abstract approach are usually free of the mentioned drawbacks, as they only generate Tcl code and are entirely written in languages more suitable for implementing such extra features.
+6.9. Learning difficulty
+The direct-Tcl approach might be harder to learn for less experienced engineers. There are two reasons for such a state being:
+  1. (1)
+    The Tcl language is not considered to be the most user-friendly.
+  2. (2)
+    The user must be familiar with the EDA tool commands in the case of less common tasks, for example, working with block designs or generating vendor cores using the Tcl interface instead of the IP wizard.
+The indirect-abstract approach has a lower entry point.
+In the case of the declarative subcategory, the user must only know basic formats, such as YAML or TOML, and a few build system keywords.
+In the case of the programmable subcategory, the language of choice is Python, which is commonly considered easier to learn than Tcl.
+It is noteworthy that, in contemporary contexts, the complexity associated with programming languages and EDA tools is effectively mitigated through the application of artificial intelligence (AI) techniques (Kazemitabaar et al., 2024; Pearce et al., 2023).
+7. Summary
+Table 2 summarizes characteristics of direct-Tcl and indirect-abstract approaches for hardware build systems.
+Table 2. Direct-Tcl and indirect abstract hardware build system approaches comparison summary.
+Approach direct-Tcl indirect-abstract
+Theory of
+operation
+The build system code is executed during
+the design build flow. The vuild system code
+is executed directly by the EDA tool. The
+build system code and build code reside
+in the same domain.
+The build system code is executed before
+the design build flow. A build system generates
+Tcl or shell script that is only later run by the
+EDA tool. The build system code and build
+code reside in different domains.
+Dependency
+overhead
+and its
+implications
+Fewer dependencies. Initial environment
+setup requires fewer installations. Lower
+probability of stability problems on Long
+Term Support distributions.
+More dependencies. Initial environment
+setup requires more actions. Slightly higher
+probability of stability problems on Long
+Term Support distributions.
+Internal
+code structure
+Monolithic. The code containing logic
+for EDA tools handling is placed within
+a single file. Prefers shared global
+variables which makes build logic easier
+to follow but also bug-prone.
+Modular. The code for handling particular
+EDA tools is placed within modules in
+separate files. The encapsulation makes the
+code more bug-resistant, but the build logic
+is slightly harder to follow.
+Abstraction
+level
+Low. The common abstraction layer is
+limited to primary tasks. For example, file
+faddition, library setting, generic/parameter
+setting. Intentionally exposes EDA tool
+peculiarities to the user.
+High. The common abstraction layer tries to
+cover all areas and hide the peculiarities of
+different EDA tools. Makes porting a design
+to a different EDA tool easier.
+Constraints
+handling
+Simple. As the build code is executed
+directly by the EDA tool during the build
+system code execution, there is direct
+access to all constraint commands.
+Optional constraints addition can be
+handled with a simple if statement.
+Depends on the subcategory. In the case of
+the declarative subcategory, some less common
+constraint scenarios are more complex to
+achieve and require indirection layers. In the
+case of the programmable subcategory,
+constraints handling requires only a small
+amount of wrapper code.
+Handling EDA
+tools without
+Tcl interface
+Straightforward. These kinds of tools are
+directly called using the exec command
+or catch-eval-exec sequence if error
+handling is required. Build rerun requires
+build system code rerun.
+Slightly more complex. EDA tools without
+the Tcl interface are handled by generating
+Makefile or shell scripts. The generated build
+code scripts can be easily reused without
+rerunning the build system code.
+Arbitrary
+external
+program
+calling
+Straightforward. User simply calls Tcl
+exec command if error catching is not
+required or catch-eval-exec sequence if
+error catching is required. Programs can
+be easily called between any EDA tool
+flow stages.
+Depends on the core description file format.
+In the case of markup or data serialization
+languages, calling external programs is
+complex. In the case of, for example, Python,
+calling external programs is simple, although
+it still requires slightly more code than in Tcl.
+Extra features
+support
+Only HBS supports automatic testbench
+detection, parallel testbench running, and
+dependency graph generation. No tool
+natively supports distributed execution of
+compilation flows.
+Highly depends on the tool. For example,
+SiliconCompiler is the only tool natively
+supporting the distributed execution of
+compilation flows. Only SiliconCompiler,
+bender and orbit support core dependency
+graph generation.
+Extra features
+implementation
+complexity
+Higher. Extra features, for example,
+parallel testbench running, require more
+code. Moreover, the code is asynchronous
+and has a non-linear structure. The build
+system code gets fragmented.
+Lower. Build systems representing this
+approach don’t have to be implemented in
+Tcl. Most of them use Python, which
+makes implementing extra features easier
+and does not split the build system code logic.
+Learning
+difficulty
+Higher. The user must directly write Tcl.
+Moreover, the user has to be familiar with
+EDA tools commands in the case of less
+common tasks, for example, working with
+block designs.
+Lower. Markup, data serialization, and Python
+syntaxes are much simpler and easier to read
+even for newcomers.
+The taxonomical data presented in the table can serve as a strategic framework for selecting an optimal hardware build system based on specific project requirements and organizational constraints.
+For example, in the case of massive-scale ASIC projects that scale to even thousands of developers and millions of servers, the indirect-abstract approach is currently the only viable solution.
+Valid choices might be SiliconCompiler or HAMMER, as both provide the necessary abstraction layers to manage this complexity.
+For small-to-medium projects with a low entry threshold, a pure declarative, indirect-abstract hardware build system might be a good choice, such as FuseSoc.
+In the case of tight coupling with Git and CI/CD, Hog might be the best choice.
+For medium-to-large projects targeting a single-vendor ecosystem, a direct-Tcl approach (e.g., HBS) is often superior.
+The direct-Tcl approach is particularly effective for teams with deep expertise in EDA tool command interfaces, where the steeper learning curve is offset by granular control over the implementation flow.
+Beyond selection, the comparative analysis in Table 2 can provide a roadmap for developing next-generation hardware build systems.
+By identifying the inherent bottlenecks and architectural weaknesses of existing approaches, developers can proactively implement mitigation strategies.
+Understanding these trade-offs a priori allows for the design of more resilient, scalable hardware build systems that address the limitations of current direct-Tcl and indirect-abstract tools.
 Acknowledgements.